### 11 Jan 2024 v2

# Tiny Time Mixers (TTMs): Fast Pretrained Models for Enhanced Zero/Few-Shot Forecasting of Multivariate Time Series

Abstract
[…]
1 Introduction
Multivariate time series (TS) forecasting entails predicting future values for multiple interrelated time series based on their historical data. This field has advanced significantly, applying statistical and machine learning (ML) methods Hyndman and Athanasopoulos (2021) across domains like weather, traffic, retail, and energy. In general, each time series represents a variable or channel111“Channel” refers to the individual time series in multivariate data (i.e., a multivariate TS is a multi-channel signal).. In certain applications, non-forecasting variables, categorized as controllable and uncontrollable external factors, impact the variables to forecast. We term these non-forecasting variables as exogenous, and the variables requiring forecast as target variables.
Related Work: Recent advances in multivariate forecasting have been marked by the advent of transformer-based Vaswani et al. (2017) approaches, exemplified by models like PatchTST Nie et al. (2022), Autoformer Wu et al. (2021), Informer Zhou et al. (2021), and FEDFormer Zhou et al. (2022). These models have demonstrated notable improvements over traditional statistical and ML methods. Furthermore, architectures based on MLP-Mixer Tolstikhin et al. (2021), such as TSMixer Ekambaram et al. (2023), have emerged as efficient transformer alternatives, boasting 2-3X reduced compute and memory requirements with no accuracy compromise compared to their transformer counterparts. However, none of these advanced approaches have successfully demonstrated the ability to create general pretrained models that can successfully transfer the learning to unseen target TS dataset, in a similar way as popularly witnessed in NLP and vision tasks. This is very challenging in the TS domain due to the diverse nature of the datasets across applications and the limited public availability of pretraining TS data. There are existing self-supervised pretraining TS approaches using masked modeling and contrastive learning techniques such as SimMTM Dong et al. (2023) and TF-C Zhang et al. (2022) that offer transfer learning between two datasets when carefully selected based on the dataset properties. However, they fail to provide universal transfer learning capabilities across datasets. Consequently, there has been a recent growing trend to employ pretrained large language models (LLMs) for TS forecasting, treating it as a cross-domain transfer learning task. These universal cross-transfer approaches, specifically recent works such as LLMTime Gruver et al. (2023) and GPT4TS Zhou et al. (2023) yield promising results in few/zero-shot forecasting approaches. These models are bootstrapped from GPT-2/3 or LLAMA-2 with suitable tokenization strategies to adapt to time-series domains.
However, these LLM based TS approaches do not explicitly handle channel correlations and exogenous support in the context of multivariate forecasting. Moreover, these large models, with billions of parameters, demand significant computational resources and runtime. Hence, in this paper, we focus on building pretrained models from scratch solely using TS data. Unlike language, which has abundant public pretraining data in terabytes, time-series data is relatively scarce, very diverse and publicly limited. Its scarcity leads to overfitting when pretraining “large” models solely on time-series data. This prompts a question: Can smaller models pretrained purely on limited public diverse TS datasets give better zero/few-shot forecasting accuracy? Surprisingly, the answer is yes! Toward this, we propose Multi-level Tiny Time Mixers (TTM), a significantly smaller model ($\leq$1M parameters) based on the lightweight TSMixer architecture, exclusively trained on diverse TS corpora for effective zero/few-shot multivariate TS forecasting via transfer learning.
In particular, TTM is pretrained using multiple public datasets ($\sim$ 240M samples) from the Monash data repository [^2] Accessible at https://forecastingdata.org/ Godahewa et al. (2021)). Note that the datasets exhibit considerable diversity in terms of characteristics, such as the different domains, temporal resolution [^3] Resolution refers to the sampling rate of the input time series (e.g., hourly, 10 minutes, 15 minutes, etc.) (spanning from second to daily), lengths, and number of channels. Pretraining on such heterogeneous datasets cannot be handled directly by TSMixer or existing state-of-the-art (SOTA) models. Hence, TTM proposes the following enhancements to the TSMixer architecture: (i) Adaptive Patching across layers, considering the varied suitability of patch lengths for different datasets, (ii) Dataset Augmentation via Downsampling to increase coverage and samples across different resolutions, (iii) Resolution Prefix Tuning to explicitly embed resolution information in the first patch, facilitating resolution-conditioned modeling, particularly beneficial in scenarios with short history lengths. Moreover, our approach leverages multi-level modeling, where TTMs are first pretrained in a channel-independent way and then seamlessly integrate channel mixing during fine-tuning to model target data-specific channel-correlations and exogenous infusion
Below, we outline the paper’s key contributions:
  • Amidst the prevalence of large pretrained models demanding significant compute and training time, our work is the first to showcase the efficacy of building Fast and Tiny Pretrained models ($\leq$1M parameters) exclusively trained on Public TS datasets in a flash of just 4-8 hours (using 6 A100 GPUs). TTM successfully demonstrates transfer learning to diverse, unseen target datasets for zero/few-shot forecasting, addressing the data scarcity issues prevalent in time series.
  • Pretraining on heterogeneous multi-resolution datasets cannot be handled effectively by TSMixer or other SOTA models. Hence, we propose various architectural and training enhancements, such as adaptive patching, data augmentation via downsampling, and (an optional) resolution prefix tuning for robust pretraining.
  • TTM employs a multi-level modeling strategy to explicitly model channel-correlations, and incorporates exogenous signals – a crucial capability lacking in LLMs-based TS approaches.
  • With extensive evaluation on 11 datasets, TTM shows significant accuracy gains over popular benchmarks (12-38% in few/zero-shot forecasting). It also drastically reduces the compute needs as compared to LLM-TS methods, with a 14X cut in learnable parameters, 106X less total parameters, and substantial reductions in finetuning (65X), inference time (54X), and memory usage (27X).
  • The zero-shot results of TTM often surpass the few-shot results of many SOTA approaches, highlighting the effectiveness of our approach.
Figure 1: Overview of Multilevel Tiny Time Mixers (TTM): (a) Refer to Section 2 and 3, (b) Refer to Section 3.1, (c) Refer to Section 3.2
2 TTM Components
Let $\bm{X}\in\mathbb{R}^{c\times sl}$ be a multivariate time series of length $sl$ and $c$ number of channels. The forecasting task can be formally defined as predicting the future values $\bm{Y}\in\mathbb{R}^{c^{\prime}\times fl}$ given the history $\bm{X}$. Here, $fl$ denotes the forecast horizon and $c^{\prime}$ denotes the forecast channels, where $c^{\prime}\leq c$. The predictions from the model are denoted by $\hat{\bm{Y}}\in\mathbb{R}^{c^{\prime}\times fl}$. In a general multivariate forecasting task, each channel or variable falls into one of the following categories: (a) Target variables (mandatory):corresponding to the channels ($c^{\prime}$) for which forecasts are required, (b) Exogenous variables (optional):encompassing (i) uncontrolled variables that may influence the forecasts and assumed to be known or estimated for the forecast period (e.g. weather), and (ii) control variables whose future values during the forecast horizon can be manipulated to govern the behavior of the target variables. (e.g. discount in sales forecasting, operator controls in industrial applications). In TTM, uncontrolled and control variables are treated similarly, as both are considered available during forecasting.
2.1 Multi-level Modeling
TTM follows a multi-level architecture consisting of four key components (see Figure 1(a)): (1) The TTM Backbone is assembled using building blocks derived from the efficient TSMixer architecture Ekambaram et al. (2023). TSMixer is based on simple MLP blocks that enable mixing of features within patches, across patches and channels, surpassing existing transformer-based TS approaches with minimal computational requirements. Since TSMixer is not targeted to handle multi-resolution data, we introduce various novel enhancements to it as explained later. (2) TTM Decoderfollows the same backbone architecture but is considerably smaller in size, approximately 10-20% of the size of the backbone, (3) Forecast Headconsists of a linear head designed to produce the forecast output, and (4) Optional Exogenous Mixer serves to fuse exogenous data into the model’s forecasting process. This multi-level model refactoring is required to dynamically change the working behavior of various components based on the workflow type, as explained in Section 3. In addition to the above primary components, we also have a preprocessing component as explained next.
2.2 Pre-processing
As shown in Figure 1(a) with colorless blocks, the historical time series $\bm{X}$ is first normalized per instance to have zero mean and unit standard deviation for each channel dimension, to tackle any possible distribution shifts Nie et al. (2022); Ekambaram et al. (2023). This process is reversed at the end before computing the loss. The normalized data $\overline{\bm{X}}$ is subsequently patched $\bm{X}_{p}\in\mathbb{R}^{c\times n\times pl}$ into $n$ non-overlapping windows, each of length $pl$ and then, passed to the TTM backbone. Patching, as introduced in Nie et al. (2022), has proven to be highly valuable for forecasting. Its effectiveness lies in preserving local semantic information, accommodating longer history, and reducing computation.
3 TTM Workflows
TTM works in two stages: pretrain and finetune (Figure 1).
3.1 Pretraining Workflow
In the pretraining stage, we train the model on a large collection of public datasets from the Monash data repositoryGodahewa et al. (2021). Since the primary focus of TTM is forecasting, pretraining is modeled with a direct forecasting objective. TTM is first pretrained in a univariate fashion with independent channels on all the existing datasets. Due to varied channel counts in pretraining datasets, modeling multivariate correlations is not feasible here; it is addressed later during fine-tuning. Multivariate pretraining datasets are initially transformed into independent univariate time series $(\bm{X}_{1},\cdots,\bm{X}_{N})\in\mathbb{R}^{c(=1)\times sl}$. These are pre-processed (Section 2.2), and subsequently fed into the TTM backbone for multi-resolution pretraining. The output of the backbone $\bm{X}_{h}^{L}\in\mathbb{R}^{(c=1)\times n\times hf}$ is passed through the decoder and forecast head to produce the forecast $\hat{\bm{Y}}\in\mathbb{R}^{(c=1)\times fl}$ which is then reverse-normalized to bring back to the original scale. We pretrain the TTM with mean squared error (MSE) loss function calculated over the forecast horizon: $\mathcal{L}=||\bm{Y}-\hat{\bm{Y}}||_{2}^{2}$. Thus for a given input context length $sl$ and forecast length $fl$, we get a pretrained model capturing the common temporal forecasting dynamics and seasonal patterns as observed in the overall pretraining data.
3.1.1 Multi-Resolution Pretraining via TTM Backbone
The majority of the pretraining happens in the TTM backbone. The primary challenge with the proposed pretraining technique is that the pretraining data is diverse and has multiple resolutions. There are two main options for pretraining: conducting separate pretraining for each resolution type or pretraining using all resolution data collectively. While it’s common to train a model per resolution type to overcome challenges in learning diverse seasonal patterns, this leads to diminished training data for each resolution due to limited data availability. Consequently, this motivated the exploration of pretraining a single model using datasets from all resolutions. To achieve this, we propose the following 3 enhancements.
Adaptive Patching: The TTM backbone is crafted with an adaptive patching architecture where different layers of the backbone operate at varying patch lengths and numbers of patches. Since each dataset in the pretraining corpora may perform optimally at a specific patch length, this approach greatly aids in generalization when diverse datasets with different resolutions are introduced. As shown in Figure 1(b), the patched data $\bm{X}_{p}\in\mathbb{R}^{c\times n\times pl}$ is passed through a embedding layer to project it to the patch hidden dimension, $\bm{X}_{h}\in\mathbb{R}^{c\times n\times hf}$. Optionally, if the resolution prefix tuning module is activated (as explained later), the resolution prefix is concatenated with $\bm{X}_{h}$. For notational simplicity, we denote the concatenated tensor with $\bm{X}_{h}$ as well.
The TTM backbone consists of $L$ levels, each comprising $M$ TTM blocks with identical patch configurations. Each TTM block is comprised of a patch partition model, a vanilla TSMixer block, and a patch merging block. A TTM block in the $i$-th level, $i=1,\ldots,L$, receives the processed data $\bm{X}_{h}^{(i-1)}\in\mathbb{R}^{c\times n\times hf}$ from the earlier block, and performs the following operations on it. Patch Partition block at every level $i$ increases the number of patches by a factor of $K_{i}$ and reduces the patch dimension size by the same factor by reshaping $\bm{X}_{h}^{(i-1)}\in\mathbb{R}^{c\times n\times hf}\text{to }\bm{X}_{h}^{i}\in\mathbb{R}^{c\times(n\cdot K_{i})\times(hf/K_{i})}$, where $K_{i}=2^{(L-i)}$. Note that, we set $hf=2^{m}$ for some integer $m\geq L$. Then, TSMixer is applied to the adapted data $\bm{X}_{h}^{i}$. Finally, the output from TSMixer is again reshaped to its original shape (i.e., $\mathbb{R}^{c\times n\times hf}$) in the patch merging block. Note that, as we go deeper into the network, the number of patches decreases while the patch dimension size increases leading to adaptive patching which helps in better generalization as we pretrain with multiple datasets together. This idea of adaptive patching is popular and very successful in the vision domain (E.g. Swin transformers Liu et al. (2021) and we are the first to port it successfully to the time-series domain to resolve multi-resolution issues in pretraining with diverse TS datasets. Figure 1(b) shows the TTM backbone for $L=3$ and $M=2$.
Data Augmentation via Downsampling: A significant challenge in TS pretraining datasets is the scarcity of public datasets at specific resolutions. To overcome this, we employ a downsampling technique for high-resolution datasets, generating multiple datasets at lower resolutions. For example, from a one-second resolution dataset, we derive datasets at minute and hour resolutions. Importantly, the original high-resolution dataset remains within the pool of pretraining datasets. This methodology significantly augments the number of datasets for each resolution which greatly improves the model performance (Section 4.5).
Resolution Prefix Tuning: This technique explicitly learns and incorporates a new patch embedding as a prefix into the input data based on the input resolution type (see Figure 1(b)). Similar to the concept of prefix tuning Li and Liang (2021), this approach provides an explicit signal to the model about the resolution type for resolution-conditioned modeling. First, we map every resolution to a unique integer, which is then passed through an embedding layer to project it to the hidden dimension, $hf$. Subsequently, we expand the embedding across all channels to have a representation of shape $c\times 1\times hf$. This module is optional for the TTM backbone, particularly beneficial when the context length ($sl$) is short. In those scenarios, automatically detecting the resolution becomes a challenge for the model. Hence, by explicitly fusing the resolution information as a prefix, we can enhance the model’s ability to learn effectively across resolutions.
3.2 Fine-tuning Workflow
In the fine-tuning workflow, we deal with data from the target domain that has no overlap with the pretraining datasets. We have three options here: (a) In Zero-shot forecasting, we directly use the pretrained model to evaluate on the test part of the target data, (b) In Few-shot forecasting, we utilize only a tiny portion (5-10%) of the train part of the target data to quickly update the pretrained weights of the decoder and head, and subsequently, evaluate it on the test part, (c) In Full-shot forecasting, we finetune the pretrained weights of the decoder and head on the entire train part of the target data, and then, evaluate on the test part.
The backbone is completely frozen during fine-tuning, and still operates in a channel-independent univariate fashion. However, the TTM decoder can be fine-tuned via channel-mixing (for multivariate) or a channel-independent (for univariate) way based on the nature of the target data. If pure multivariate modeling is needed, then the channel-mixer block in all the TSMixer components (see Figure 1(b)) in the decoder gets enabled to explicitly capture the channel correlation between the channels. The forecast head and reverse normalization perform similar operations as in the pretraining stage. The fine-tuning also optimizes the forecasting objective with MSE loss. This thoughtful multi-level design choice ensures our backbone excels in channel-independent pretraining, enabling effective temporal correlation modeling across diverse datasets. Simultaneously, the decoder handles target-data-specific tasks like channel-correlation modeling and fine-tuning. In addition, if the target data has exogenous variables, then an exogenous mixer block is applied to the actual forecasts as explained next.
Exogenous Mixer Block: As described in Section 2, the future values of the exogenous channels are known in advance. Let the forecast from the forecast head be $\hat{\bm{Y}}\in\mathbb{R}^{c\times fl}$. Let the channels $\bm{x}_{0},\cdots,\bm{x}_{N_{1}}$ denote the target variables and $\bm{x}_{N_{1}+1},\cdots,\bm{x}_{N_{1}+N_{2}}$ denote all exogenous variables with their future values known. First, we replace the forecast values for the exogenous channels with the true future values ($\bm{Y}$) and transpose it: $\hat{\bm{Y}}_{e}=\left[\hat{\bm{y}}_{0},\cdots,\hat{\bm{y}}_{N_{1}},\bm{y}_{N_{1}+1},\cdots,\bm{y}_{N_{1}+N_{2}}\right]\in\mathbb{R}^{fl\times c}$. Next, to learn inter-channel lagged correlations, we patch $\hat{\bm{Y}}_{e}$ into a series of overlapped windows (i.e., patching with stride$=1$) to create a new tensor: $\hat{\bm{Y}}_{e,p}\in\mathbb{R}^{fl\times\Delta\times c}$, where $\Delta=2\cdot l+1$ with $l$ being the context length to incorporate on either side of a time point [^4] This needs padding $\hat{\bm{Y}}_{e}$ with zeros of length $l$ on both sides.. Subsequently, we pass $\hat{\bm{Y}}_{e,p}$ through a vanilla TSMixer block with channel mixing enabled. Thus, the lagged dependency of the forecasts for the target channels on the exogenous channels is seamlessly learned. Finally, we attach a linear head to produce the forecasts for the target channels which is then reshaped as $\hat{\bm{Y}}\in\mathbb{R}^{N_{1}\times fl}$. Figure 1(c) depicts this procedure.
Zero-shot Few-shot (5%) $fl$ TTM TTM GPT4TS PatchTST TSMixer TimesNet Dlinear FEDFormer Autoformer ETTH1 96 0.365 0.366 0.543 0.557 0.554 0.892 0.547 0.593 0.681 192 0.393 0.391 0.748 0.711 0.673 0.940 0.720 0.652 0.725 336 0.415 0.421 0.754 0.816 0.678 0.945 0.984 0.731 0.761 720 0.538 - - - - - - - - ETTH2 96 0.285 0.282 0.376 0.401 0.348 0.409 0.442 0.390 0.428 192 0.341 0.338 0.418 0.452 0.419 0.483 0.617 0.457 0.496 336 0.383 0.383 0.408 0.464 0.389 0.499 1.424 0.477 0.486 720 0.441 - - - - - - - - ETTM1 96 0.413 0.359 0.386 0.399 0.361 0.606 0.332 0.628 0.726 192 0.476 0.402 0.440 0.441 0.411 0.681 0.358 0.666 0.750 336 0.553 0.424 0.485 0.499 0.467 0.786 0.402 0.807 0.851 720 0.737 0.575 0.577 0.767 0.677 0.796 0.511 0.822 0.857 ETTM2 96 0.187 0.174 0.199 0.206 0.200 0.220 0.236 0.229 0.232 192 0.261 0.240 0.256 0.264 0.265 0.311 0.306 0.394 0.291 336 0.323 0.299 0.318 0.334 0.314 0.338 0.380 0.378 0.478 720 0.436 0.407 0.46 0.454 0.410 0.509 0.674 0.523 0.553 Weather 96 0.154 0.152 0.175 0.171 0.188 0.207 0.184 0.229 0.227 192 0.203 0.198 0.227 0.230 0.234 0.272 0.228 0.265 0.278 336 0.256 0.250 0.286 0.294 0.287 0.313 0.279 0.353 0.351 720 0.329 0.326 0.366 0.384 0.365 0.400 0.364 0.391 0.387 Electricity 96 0.169 0.142 0.143 0.145 0.147 0.315 0.150 0.235 0.297 192 0.196 0.162 0.159 0.163 0.172 0.318 0.163 0.247 0.308 336 0.209 0.184 0.179 0.175 0.190 0.340 0.175 0.267 0.354 720 0.264 0.228 0.233 0.219 0.280 0.635 0.219 0.318 0.426 Traffic 96 0.518 0.401 0.419 0.404 0.408 0.854 0.427 0.670 0.795 192 0.548 0.425 0.434 0.412 0.421 0.894 0.447 0.653 0.837 336 0.55 0.437 0.449 0.439 0.477 0.853 0.478 0.707 0.867 720 0.605 - - - - - - - - TTM Zero-shot % IMP 1% 3% 2% 31% 6% 25% 32% TTM 5% Few-shot % IMP 12% 14% 12% 38% 17% 32% 38% Table 1: Zero-shot and Few-shot 5% performance of TTM and all SOTA models on D1 datasets for varying forecast lengths ($fl$). Bold and underscore denote the best and the second best. For longer $fl$ (720), some datasets fail to generate fine-tuning samples with just 5% of the data, denoted as ‘-’. The last two rows show relative improvement (IMP) of TTM zero-shot and few-shot w.r.t. SOTA few-shot results.
Metric Data GPT4TS TTM (Ours) IMP (X) FT NPARAMS (M) Weather 3.2 0.29 (14X) Electricity 3.2 0.29 Traffic 5.5 0.29 TL NPARAMS (M) Weather 84.3 0.8 (106X) Electricity 84.3 0.8 Traffic 86.6 0.8 EPOCH TIME (s) Weather 37 1 (65X) Electricity 241 4 Traffic 782 8 MAX MEMORY (GB) Weather 21.9 1.06 (27X) Electricity 39.7 2.02 Traffic 53.7 1.34 TEST TIME (s) Weather 52.53 1.66 (54X) Electricity 353.4 8.07 Traffic 1163.9 13.5 Table 2: Computational Improvement in few-shot 10% for $fl$=96 using 1 A100 GPU. nX denotes n times average improvement across datasets (Avg. IMP). 10% 25% 50% 75% 100% IMP TTM 0.422 0.421 0.413 0.402 0.398 - SimMTM 0.591 0.535 0.491 0.466 0.415 17% Ti-MAE 0.660 0.594 0.55 0.475 0.466 24% TST 0.783 0.641 0.525 0.516 0.469 28% LaST 0.645 0.610 0.540 0.479 0.443 23% TF-C 0.799 0.736 0.731 0.697 0.635 43% CoST 0.784 0.624 0.540 0.494 0.428 26% TS2Vec 0.655 0.632 0.599 0.577 0.517 31%
Figure 2: Cross transfer learning improvement (IMP) for self-supervised pretraining methods in various few-shot settings (10%,25%,50%,75%,100%).
4 Experiments and Results
4.1 Experimental Setting
Datasets: Pretraining employs a subset of the Monash data hub Godahewa et al. (2021) of size $\sim 240$M samples. We specifically exclude datasets that consist of yearly, monthly, and daily series as they do not possess sufficient length for the long-term forecasting task. Moreover, we remove all the datasets that we utilize for evaluation. This includes weather, electricity, and traffic datasets. For zero/few/full-shot evaluation we consider seven datasets (D1): ETTH1, ETTH2, ETTM1, ETTM2, Weather, Electricity (ECL), Traffic as popularly used in most prior SOTA works Zhou et al. (2021); Nie et al. (2022). Since these datasets do not contain any exogenous variables nor exhibit cross-channel correlation benefits, we incorporate four other datasets (D2) for separately validating the efficacy of the decoder channel mixing and exogenous mixer module: bike sharing (BS) Fanaee-T (2013), carbon capture plant (CC) Jablonka et al. (2023), and two more datasets from Business-IT observability domain Biz (2023): Application (APP) and Service (SER). Details are provided in the Appendix.
SOTA Bechmarks: We benchmark TTM model with the latest publicly available SOTA models for multivariate forecasting categorized as follows: [label=()] 1. LLM-based TS pretrained models:GPT4TS Zhou et al. (2023), LLMTime Gruver et al. (2023) 2. Self-supervised pretrained models: SimMTM Dong et al. (2023),Ti-MAE Li et al. (2023), TST Zerveas et al. (2021), LaST Wang et al. (2022), TF-C Zhang et al. (2022), CoST Woo et al. (2022) and Ts2Vec Yue et al. (2022) 3. TS transformer models:PatchTST Nie et al. (2022), FEDFormer Zhou et al. (2022), Autoformer Wu et al. (2021) 4. Other SOTA models:TSMixer Ekambaram et al. (2023), DLinear Zeng et al. (2022) and TimesNet Wu et al. (2022)
Zero-shot Few-shot (10%) Data TTM (Ours) TTM (Ours) GPT4TS PatchTST TSMixer TimesNet Dlinear FEDFormer Autoformer ETTH1 0.428 0.422 0.590 0.633 0.578 0.869 0.691 0.639 0.702 ETTH2 0.362 0.360 0.397 0.415 0.370 0.479 0.605 0.466 0.488 ETTM1 0.545 0.458 0.464 0.501 0.441 0.677 0.411 0.722 0.802 ETTM2 0.302 0.278 0.293 0.296 0.282 0.320 0.316 0.463 1.342 Electricity 0.210 0.180 0.176 0.176 0.180 0.323 0.180 0.346 0.431 Traffic 0.555 0.428 0.440 0.430 0.429 0.951 0.446 0.663 0.749 Weather 0.236 0.227 0.238 0.242 0.236 0.279 0.241 0.284 0.300 TTM Zero-shot improvements -4% -2% -7% 27% 2% 27% 39% TTM Few-shot (10%) improvements 7% 9% 4% 34% 13% 34% 45% Table 3: Zero-shot and Few-shot 10% performance of TTM and all SOTA models on seven datasets averaged across 4 forecast lengths. The numbers with bold and underscore denote the best the second best results respectively. The full table is in Appendix.
Data $fl$ Zero-shot LLMTime TTM Traffic 96 0.340 0.195 192 0.526 0.367 Weather 96 0.107 0.031 192 0.062 0.067 Electricity 96 0.609 0.238 192 0.960 0.488 ETTM2 96 0.167 0.164 192 0.198 0.228 TTM IMP (%) 29% Table 4: Comparison between TTM and LLMTime in zero-shot setting.
TTM Model Details: By default, the following data and model configuration is used for pretraining: input sequence length $sl$ = 512, patch length $pl$ = 64, stride $s$ = 64, number of patches $n=512/64=8$, forecast length $fl\in\{96,192,336,720\}$, number of levels $L$ = 6, number of TTM blocks per level $M$ = 2, batch size $b$ = 3000, number of epochs $ep$ = 25, and dropout $do$ = 0.2. TSMixer specific hyperparameters include feature scaler $fs$ = 3, hidden feature size $hf$ = $fs*pl=192$, expansion feature size $ef=hf*2=384$. Please note that, $hf$ and $n$ will change across TTM blocks based on the adaptive patching strategy, and the decoder works in channel-independent mode during pretraining. Resolution prefix tuning is disabled by default and enabled only for shorter context length (as explained in Table. 9). Training is performed in a distributed fashion with 50 CPUs and 6 NVIDIA A100 GPUs with 40 GB GPU memory. We build a pretrained model using Monash data for each specified $fl$. Since pretrained models are very small, they can be trained quickly in a few hours (4-8 hrs based on $fl$ length), as opposed to several days/weeks in standard approaches. For few/full-shot fine-tuning, all the above pretraining parameters remain the same except the following. Based on the target data characteristics, the decoder channel-mixing and/or exogenous mixer block is enabled/disabled (both enabled for D2 datasets). In addition, head dropout changes to 0.7 for smaller ETT datasets and 0.2 otherwise and batch size ranges between 8-64. All the hyperparameters are adjusted based on the validation performance, and the final test results are reported in the paper.
Data BS CC APP SER % IMP
TTM-CM 0.582 0.250 0.042 0.114
TTM-Zero-shot 0.992 0.263 0.183 0.238 44%
TTM 0.635 0.261 0.073 0.143 18%
PatchTST 0.735 0.267 0.060 0.119 15%
TSMixer-CC 0.651 0.284 0.053 0.136 15%
TSMixer-CM 0.716 0.303 0.069 0.118 20%
TSMixer 0.664 0.267 0.066 0.134 17%
GPT4TS 0.645 0.254 0.075 0.135 18%
Table 5: Effect of Decoder Mixing and Exog Fusion. Results are reported using $(sl,fl)$ with values of $(512,96)$ for BS and $(96,24)$ for other datasets.
4.2 TTM’s Zero/Few-shot Performance
Table 1 compares the performance of TTM model in zero-shot and few-shot (5%) settings. Baseline results are reported from  Zhou et al. (2023) as we use the same few-shot data filtering strategy as followed in that paper. Note that for zero-shot performance, the TTM model pretrained on the Monash data is directly evaluated on the test set for every dataset. We report the test MSE for multiple forecast horizons varying from 96 to 720. The 5% few-shot TTM outperforms the SOTAs in most of the cases with significant accuracy gains (12-38%). An even more impressive observation is that the TTM in a zero-shot setting is also able to outperform most of the SOTAs which are trained on 5% of the target data. This observation establishes the generalization ability of the pretrained TTM model on the target datasets. Likewise, Table 3 shows the 10% few-shot performance of the TTM, where we outperform all the existing SOTAs by 4-45% accuracy gains. In addition, TTM zero-shot also beats many SOTAs (not all) with 10% training highlighting the effectiveness of our approach.
Additionally, we conduct a comparison between TTM and the LLMTime model which is explicitly designed for zero-shot setting. Since these models are based on LLaMA and are massive in size, the authors used only the last window of the standard test-set for faster evaluation, as opposed to using all the windows in the full test-set . Hence, we compare them separately in Table 4 based on the same datasets, $fl$, and test-set as reported in  Gruver et al. (2023). In this evaluation, we outperform LLMTime by a substantial margin of 29%. In addition, there are alternative pretraining approaches like masked modeling and contrastive learning techniques that may not offer universal transfer learning capabilities across all datasets (like TTM), but they excel in enabling cross-transfer learning between two datasets when carefully selected. Table 2 illustrates the cross-transfer learning from ETTH2 to ETTH1 for these models in various few-shot settings (as reported in  Dong et al. (2023)). Notably, TTM, with no specific cross-data selection, outperforms all popular SOTAs, including the latest SimMTM, by 17-43%. Thus, TTM significantly beats all the existing SOTAs including the recent popular transfer learning benchmarks based on LLMs. The notable part here is that we get these substantial accuracy gains with significantly reduced compute/runtime requirements as explained next.
Data Zero-shot Few-shot (5%)) Few-shot (10%))
5 Epochs 50 Epochs 5 Epochs 50 Epochs
RI PT RI PT RI PT RI PT RI PT
ETTH1 0.735 0.428 0.687 0.389 0.629 0.393 0.64 0.419 0.552 0.422
ETTH2 0.400 0.362 0.378 0.334 0.379 0.334 0.371 0.357 0.370 0.360
ETTM1 0.722 0.545 0.474 0.483 0.448 0.440 0.421 0.458 0.403 0.458
ETTM2 0.353 0.302 0.296 0.274 0.290 0.280 0.279 0.271 0.274 0.278
Electricity 0.890 0.210 0.356 0.188 0.199 0.179 0.220 0.186 0.178 0.180
Traffic 1.453 0.555 0.593 0.454 0.441 0.421 0.486 0.446 0.430 0.428
Weather 0.321 0.236 0.267 0.233 0.264 0.231 0.240 0.229 0.238 0.227
IMP 36% 21% 12% 9% 2%
Table 6: Effect of pretraining (PT) over Randomly Initialized (RI) weights, averaged across 4 values of $fl$. Full table in Appendix.
Data Zero-shot Few-shot (10%) Vanilla Adaptive Patching Vanilla Adaptive Patching ETTH1 0.501 0.428 0.478 0.422 ETTH2 0.375 0.362 0.372 0.360 ETTM1 0.577 0.545 0.463 0.458 ETTM2 0.310 0.302 0.272 0.278 Electricity 0.215 0.210 0.178 0.180 Traffic 0.556 0.555 0.431 0.428 Weather 0.239 0.236 0.230 0.227 IMP (%) 4% 2% Table 7: Effect of adaptive Patching, averaged across 4 values of $fl$.
Data Zero-shot Original Data With Downsample ETTH1 0.603 0.379 ETTH2 0.351 0.313 ETTM1 0.852 0.444 ETTM2 0.261 0.224 Electricity 0.424 0.182 Traffic 0.898 0.533 Weather 0.186 0.178 IMP (%) 30% Table 8: Effect of downsampling, averaged across $fl$ 96 and 192.
Data $HL$ = 96 $HL$ = 512 NP P NP P ETTH1 0.373 0.358 0.365 0.360 ETTH2 0.180 0.179 0.285 0.280 ETTM1 0.559 0.387 0.413 0.384 ETTM2 0.127 0.108 0.187 0.194 Weather 0.103 0.103 0.154 0.160 Electricity 0.208 0.201 0.169 0.175 Traffic 0.754 0.740 0.518 0.518 IMP (%) 8% 0% Table 9: Resolution Prefix Tuning: With (P) vs Without(NP)
4.3 Computational Benefits of TTM
Table  2 compares the computational benefits of TTM with GPT4TS, the most popular LLM based TS model and the current best SOTA in few-shot setting. We run both models in a single A100 GPU with matching batch sizes for a fair comparison. Multi-GPU usage is deliberately avoided in this experiment to eliminate IPC overheads and ensure precise metric calculations. Since GPT4TS is based on GPT-2, it consumes a huge model footprint and exhibits slow execution as compared to TTM. In specific, TTM achieves 14X cut in Finetune (FT) parameters (NPARAMS) and a remarkable 106X reduction in Total (TL) parameters. This reduction in model footprint further leads to a significant reduction in fine-tuning EPOCH TIME by 65X, MAX MEMORY usage by 27X, and total inference time on the entire test data (TEST TIME) by 54X. The computational advantages also apply over other LLM-based time series models like LLMTime, built on LLaMA, which is larger than GPT-2 (utilized in GPT4TS). In fact, LLMTime used a very small test set on these datasets as opposed to the standard test set to overcome this slow-execution behavior.
4.4 TTM’s effectiveness in cross-channel and exogenous/control modeling
Since the datasets ($D1$) used in previous experiments do not have exogenous variables, we evaluate the effectiveness of TTM on 4 other datasets ($D2$, as explained in  4.1) to quantify its benefits. Table 5 shows the performance of the Monash-pretrained TTM model fine-tuned on the target data with exogenous mixer module and decoder channel-mixing enabled (TTM-CM). We compare the full-shot performance of TTM-CM with the pretrained TTM used in zero-shot setting (TTM-Zero-shot), and plain TTM (TTM) and other primary SOTAs (PatchTST, TSMixer variants, and GPT4TS) trained from scratch. Remaining SOTAs are not reported in Table 5 considering their inferior performance and space constraints. Specifically, we compare with TSMixer with channel-mixing enabled (TSMixer-CM) and TSMixer with Cross-Channel reconciliation head (TSMixer-CC) Ekambaram et al. (2023) as they are latest SOTAs in channel-correlation modeling. We can see that TTM-CM outperforms all the competitive models with a significant margin (15-44%), thus, demonstrating that the power of TTM in capturing inter-channel correlations that are prevalent in purely multivariate datasets.
4.5 Ablation Studies
Effect of Pretraining: Table 6 illustrates the advantages of the proposed pretraining (PT) approach in comparison to randomly initialized (RI) model weights. In the case of zero-shot, the pretrained TTM (PT) exhibits 36% improvement over RI. This outcome is expected, as random weights are directly employed for forecasting with RI. For the 5% few-shot scenario, PT achieves noteworthy improvements of 21% and 12% over RI when fine-tuned for 5 and 50 epochs, respectively. This underscores the utility of pretrained weights in facilitating quick learning with a limited amount of target data. Even in the case of 10% few-shot, PT continues to demonstrate an improved performance of 9% for 5 epochs and 2% for 50 epochs. Thus, the performance impact of leveraging pretrained weights greatly increases as the size of the training data and the available time-to-finetune reduces.
Effect of Adaptive Patching: The Adaptive patching presented in Section 3.1.1 is a key component of the TTM backbone. Table 7 demonstrates the relative improvements in MSE when we employ adaptive patching with respect to a vanilla TTM backbone without this module. We can see that adaptive patching adds an extra 4% improvement during zero-shot and 2% improvement for 10% few-shot. MSE improvements are observed in almost all datasets which justifies the effectiveness of the adaptive patching module when evaluated on multiple datasets of different resolutions.
Effect of Augmentation via Downsampling: Table 8 compares the zero-shot test performances between the TTM model trained on the original Monash data and the TTM trained on the augmented Monash data after downsampling as described in Section 3.1. It is evident that there is a significant improvement of 30% for the model pretrained on the augmented datasets as it adds more data coverage across various resolutions. This observation highlights the strength of the proposed TTM model to learn from diverse datasets together, and increasing the number of datasets for each resolution helps.
Effect of Resolution Prefix Tuning: Since resolution prefix tuning explicitly adds the resolution type as an extra embedding, it greatly helps when the input context length $sl$ is short where it is challenging for the model to derive the resolution type automatically. As observed in Table 9, we observe 8% improvement in zero-shot for shorter context length ($sl$ = 96) and no improvement when the context length is longer ($sl$ =512). Hence this component is optional and can be enabled when working with shorter input context length.
5 Conclusions and Future Work
Considering the diversity and limited availability of public TS pretraining data, pretraining large models for effective transfer learning poses several challenges in time series. Hence, we propose TTM, a multi-level Tiny Time Mixer model designed for efficient pretraining on limited diverse multi-resolution datasets. TTM achieves state-of-the-art results in zero/few-shot forecasting, offering significant computational efficiency and supporting cross-channel and exogenous variables — critical features lacking in existing popular methods. Going forward, we plan to extend our approach to many other downstream tasks beyond forecasting for a purely foundational approach in time series.

### 17 Jan 2024 v3

@@
-Tiny Time Mixers (TTMs): Fast Pretrained Models for Enhanced Zero/Few-Shot Forecasting of Multivariate Time Series
+Tiny Time Mixers (TTMs): Fast Pre-trained Models for Enhanced Zero/Few-Shot Forecasting of Multivariate Time Series
 Abstract
@@
 (2023), have emerged as efficient transformer alternatives, boasting 2-3X reduced compute and memory requirements with no accuracy compromise compared to their transformer counterparts.
-However, none of these advanced approaches have successfully demonstrated the ability to create general pretrained models that can successfully transfer the learning to unseen target TS dataset, in a similar way as popularly witnessed in NLP and vision tasks.
-This is very challenging in the TS domain due to the diverse nature of the datasets across applications and the limited public availability of pretraining TS data.
-There are existing self-supervised pretraining TS approaches using masked modeling and contrastive learning techniques such as SimMTM Dong et al.
+However, none of these advanced approaches have successfully demonstrated the ability to create general pre-trained models that can successfully transfer the learning to unseen target TS dataset, in a similar way as popularly witnessed in NLP and vision tasks.
+This is very challenging in the TS domain due to the diverse nature of the datasets across applications and the limited public availability of TS data for pre-training.
+There are existing self-supervised pre-training TS approaches using masked modeling and contrastive learning techniques such as SimMTM Dong et al.
 (2023) and TF-C Zhang et al.
@@
 However, they fail to provide universal transfer learning capabilities across datasets.
-Consequently, there has been a recent growing trend to employ pretrained large language models (LLMs) for TS forecasting, treating it as a cross-domain transfer learning task.
+Consequently, there has been a recent growing trend to employ pre-trained large language models (LLMs) for TS forecasting, treating it as a cross-domain transfer learning task.
 These universal cross-transfer approaches, specifically recent works such as LLMTime Gruver et al.
@@
 Moreover, these large models, with billions of parameters, demand significant computational resources and runtime.
-Hence, in this paper, we focus on building pretrained models from scratch solely using TS data.
-Unlike language, which has abundant public pretraining data in terabytes, time-series data is relatively scarce, very diverse and publicly limited.
-Its scarcity leads to overfitting when pretraining “large” models solely on time-series data.
-This prompts a question: Can smaller models pretrained purely on limited public diverse TS datasets give better zero/few-shot forecasting accuracy?
+Hence, in this paper, we focus on building pre-trained models from scratch solely using TS data.
+Unlike language, which has abundant public pre-training data in terabytes, time-series data is relatively scarce, very diverse and publicly limited.
+Its scarcity leads to overfitting when pre-training “large” models solely on time-series data.
+This prompts a question: Can smaller models pre-trained purely on limited public diverse TS datasets give better zero/few-shot forecasting accuracy?
 Surprisingly, the answer is yes!
 Toward this, we propose Multi-level Tiny Time Mixers (TTM), a significantly smaller model ($\leq$1M parameters) based on the lightweight TSMixer architecture, exclusively trained on diverse TS corpora for effective zero/few-shot multivariate TS forecasting via transfer learning.
-In particular, TTM is pretrained using multiple public datasets ($\sim$ 240M samples) from the Monash data repository [^2] Accessible at https://forecastingdata.org/ Godahewa et al.
+In particular, TTM is pre-trained using multiple public datasets ($\sim$ 244M samples) from the Monash data repository [^2] Accessible at https://forecastingdata.org/ Godahewa et al.
 (2021)).
 Note that the datasets exhibit considerable diversity in terms of characteristics, such as the different domains, temporal resolution [^3] Resolution refers to the sampling rate of the input time series (e.g., hourly, 10 minutes, 15 minutes, etc.) (spanning from second to daily), lengths, and number of channels.
-Pretraining on such heterogeneous datasets cannot be handled directly by TSMixer or existing state-of-the-art (SOTA) models.
+Pre-training on such heterogeneous datasets cannot be handled directly by TSMixer or existing state-of-the-art (SOTA) models.
 Hence, TTM proposes the following enhancements to the TSMixer architecture: (i) Adaptive Patching across layers, considering the varied suitability of patch lengths for different datasets, (ii) Dataset Augmentation via Downsampling to increase coverage and samples across different resolutions, (iii) Resolution Prefix Tuning to explicitly embed resolution information in the first patch, facilitating resolution-conditioned modeling, particularly beneficial in scenarios with short history lengths.
-Moreover, our approach leverages multi-level modeling, where TTMs are first pretrained in a channel-independent way and then seamlessly integrate channel mixing during fine-tuning to model target data-specific channel-correlations and exogenous infusion
+Moreover, our approach leverages multi-level modeling, where TTMs are first pre-trained in a channel-independent way and then seamlessly integrate channel mixing during fine-tuning to model target data-specific channel-correlations and exogenous infusion
 Below, we outline the paper’s key contributions:
-  • Amidst the prevalence of large pretrained models demanding significant compute and training time, our work is the first to showcase the efficacy of building Fast and Tiny Pretrained models ($\leq$1M parameters) exclusively trained on Public TS datasets in a flash of just 4-8 hours (using 6 A100 GPUs).
+  • Amidst the prevalence of large pre-trained models demanding significant compute and training time (in weeks), our work is the first to showcase the efficacy of building Fast and Tiny Pre-trained models ($\leq$1M parameters) exclusively trained on Public TS datasets in a flash of just few hours (4-8 hours, 6 A100 GPUs).
 TTM successfully demonstrates transfer learning to diverse, unseen target datasets for zero/few-shot forecasting, addressing the data scarcity issues prevalent in time series.
-  • Pretraining on heterogeneous multi-resolution datasets cannot be handled effectively by TSMixer or other SOTA models.
-Hence, we propose various architectural and training enhancements, such as adaptive patching, data augmentation via downsampling, and (an optional) resolution prefix tuning for robust pretraining.
+  • Pre-training on heterogeneous multi-resolution datasets cannot be handled effectively by TSMixer or other SOTA models.
+Hence, we propose various architectural and training enhancements, such as adaptive patching, data augmentation via downsampling, and (an optional) resolution prefix tuning for robust pre-training.
   • TTM employs a multi-level modeling strategy to explicitly model channel-correlations, and incorporates exogenous signals – a crucial capability lacking in LLMs-based TS approaches.
@@
 The forecasting task can be formally defined as predicting the future values $\bm{Y}\in\mathbb{R}^{c^{\prime}\times fl}$ given the history $\bm{X}$.
-Here, $fl$ denotes the forecast horizon and $c^{\prime}$ denotes the forecast channels, where $c^{\prime}\leq c$.
+Here, $fl$ denotes the forecast horizon and $c^{\prime}$ denotes number of forecast channels, where $c^{\prime}\leq c$.
 The predictions from the model are denoted by $\hat{\bm{Y}}\in\mathbb{R}^{c^{\prime}\times fl}$.
-In a general multivariate forecasting task, each channel or variable falls into one of the following categories: (a) Target variables (mandatory):corresponding to the channels ($c^{\prime}$) for which forecasts are required, (b) Exogenous variables (optional):encompassing (i) uncontrolled variables that may influence the forecasts and assumed to be known or estimated for the forecast period (e.g.
+In a general multivariate forecasting task, each channel or variable falls into one of the following categories: (a) Target variables (mandatory):corresponding to the channels for which forecasts are required, (b) Exogenous variables (optional):encompassing (i) uncontrolled variables that may influence the forecasts and assumed to be known or estimated for the forecast period (e.g.
 weather), and (ii) control variables whose future values during the forecast horizon can be manipulated to govern the behavior of the target variables.
@@
 3 TTM Workflows
-TTM works in two stages: pretrain and finetune (Figure 1).
-3.1 Pretraining Workflow
-In the pretraining stage, we train the model on a large collection of public datasets from the Monash data repositoryGodahewa et al.
+TTM works in 2 stages: pre-train and fine-tune (Figure 1(a)).
+3.1 Pre-training Workflow
+In the pre-training stage, we train the model on a large collection of public datasets from the Monash data repositoryGodahewa et al.
 (2021).
-Since the primary focus of TTM is forecasting, pretraining is modeled with a direct forecasting objective.
-TTM is first pretrained in a univariate fashion with independent channels on all the existing datasets.
-Due to varied channel counts in pretraining datasets, modeling multivariate correlations is not feasible here; it is addressed later during fine-tuning.
-Multivariate pretraining datasets are initially transformed into independent univariate time series $(\bm{X}_{1},\cdots,\bm{X}_{N})\in\mathbb{R}^{c(=1)\times sl}$.
-These are pre-processed (Section 2.2), and subsequently fed into the TTM backbone for multi-resolution pretraining.
+Since the primary focus of TTM is forecasting, pre-training is modeled with a direct forecasting objective.
+TTM is first pre-trained in a univariate fashion with independent channels on all the existing datasets.
+Due to varied channel counts in pre-training datasets, modeling multivariate correlations is not feasible here; it is addressed later during fine-tuning.
+Multivariate pre-training datasets are initially transformed into independent univariate time series $(\bm{X}_{1},\cdots,\bm{X}_{N})\in\mathbb{R}^{c(=1)\times sl}$.
+These are pre-processed (Section 2.2), and subsequently fed into the TTM backbone for multi-resolution pre-training.
 The output of the backbone $\bm{X}_{h}^{L}\in\mathbb{R}^{(c=1)\times n\times hf}$ is passed through the decoder and forecast head to produce the forecast $\hat{\bm{Y}}\in\mathbb{R}^{(c=1)\times fl}$ which is then reverse-normalized to bring back to the original scale.
-We pretrain the TTM with mean squared error (MSE) loss function calculated over the forecast horizon: $\mathcal{L}=||\bm{Y}-\hat{\bm{Y}}||_{2}^{2}$.
-Thus for a given input context length $sl$ and forecast length $fl$, we get a pretrained model capturing the common temporal forecasting dynamics and seasonal patterns as observed in the overall pretraining data.
-3.1.1 Multi-Resolution Pretraining via TTM Backbone
-The majority of the pretraining happens in the TTM backbone.
-The primary challenge with the proposed pretraining technique is that the pretraining data is diverse and has multiple resolutions.
-There are two main options for pretraining: conducting separate pretraining for each resolution type or pretraining using all resolution data collectively.
+We pre-train the TTM with mean squared error (MSE) loss function calculated over the forecast horizon: $\mathcal{L}=||\bm{Y}-\hat{\bm{Y}}||_{2}^{2}$.
+Thus for a given input context length $sl$ and forecast length $fl$, we get a pre-trained model capturing the common temporal forecasting dynamics and seasonal patterns as observed in the overall pre-training data.
+3.1.1 Multi-Resolution Pre-training via TTM Backbone
+The majority of the pre-training happens in the TTM backbone.
+The primary challenge with the proposed pre-training technique is that the pre-training data is diverse and has multiple resolutions.
+There are two main options for pre-training: conducting separate pre-training for each resolution type or pre-training using all resolution data collectively.
 While it’s common to train a model per resolution type to overcome challenges in learning diverse seasonal patterns, this leads to diminished training data for each resolution due to limited data availability.
-Consequently, this motivated the exploration of pretraining a single model using datasets from all resolutions.
+Consequently, this motivated the exploration of pre-training a single model using datasets from all resolutions.
 To achieve this, we propose the following 3 enhancements.
 Adaptive Patching: The TTM backbone is crafted with an adaptive patching architecture where different layers of the backbone operate at varying patch lengths and numbers of patches.
-Since each dataset in the pretraining corpora may perform optimally at a specific patch length, this approach greatly aids in generalization when diverse datasets with different resolutions are introduced.
+Since each dataset in the pre-training corpora may perform optimally at a specific patch length, this approach greatly aids in generalization when diverse datasets with different resolutions are introduced.
 As shown in Figure 1(b), the patched data $\bm{X}_{p}\in\mathbb{R}^{c\times n\times pl}$ is passed through a embedding layer to project it to the patch hidden dimension, $\bm{X}_{h}\in\mathbb{R}^{c\times n\times hf}$.
@@
 The TTM backbone consists of $L$ levels, each comprising $M$ TTM blocks with identical patch configurations.
-Each TTM block is comprised of a patch partition model, a vanilla TSMixer block, and a patch merging block.
-A TTM block in the $i$-th level, $i=1,\ldots,L$, receives the processed data $\bm{X}_{h}^{(i-1)}\in\mathbb{R}^{c\times n\times hf}$ from the earlier block, and performs the following operations on it.
+A TTM block in the $i$-th level, $i=1,\ldots,L$, receives the processed data $\bm{X}_{h}^{(i-1)}\in\mathbb{R}^{c\times n\times hf}$ from the earlier block.
+Each TTM block is further comprised of a patch partition model, a vanilla TSMixer block, and a patch merging block.
 Patch Partition block at every level $i$ increases the number of patches by a factor of $K_{i}$ and reduces the patch dimension size by the same factor by reshaping $\bm{X}_{h}^{(i-1)}\in\mathbb{R}^{c\times n\times hf}\text{to }\bm{X}_{h}^{i}\in\mathbb{R}^{c\times(n\cdot K_{i})\times(hf/K_{i})}$, where $K_{i}=2^{(L-i)}$.
@@
 Finally, the output from TSMixer is again reshaped to its original shape (i.e., $\mathbb{R}^{c\times n\times hf}$) in the patch merging block.
-Note that, as we go deeper into the network, the number of patches decreases while the patch dimension size increases leading to adaptive patching which helps in better generalization as we pretrain with multiple datasets together.
+Note that, as we go deeper into the network, the number of patches decreases while the patch dimension size increases leading to adaptive patching which helps in better generalization as we pre-train with multiple datasets together.
 This idea of adaptive patching is popular and very successful in the vision domain (E.g.
 Swin transformers Liu et al.
-(2021) and we are the first to port it successfully to the time-series domain to resolve multi-resolution issues in pretraining with diverse TS datasets.
+(2021) and we are the first to port it successfully to the time-series domain to resolve multi-resolution issues in pre-training with diverse TS datasets.
 Figure 1(b) shows the TTM backbone for $L=3$ and $M=2$.
-Data Augmentation via Downsampling: A significant challenge in TS pretraining datasets is the scarcity of public datasets at specific resolutions.
+Please note that adaptive patching is enabled only in the backbone and not in the decoder, which is designed to be very lightweight.
+Data Augmentation via Downsampling: A significant challenge in TS pre-training datasets is the scarcity of public datasets at specific resolutions.
 To overcome this, we employ a downsampling technique for high-resolution datasets, generating multiple datasets at lower resolutions.
 For example, from a one-second resolution dataset, we derive datasets at minute and hour resolutions.
-Importantly, the original high-resolution dataset remains within the pool of pretraining datasets.
+Note that, the original high-resolution dataset remains within the pool of pre-training datasets.
 This methodology significantly augments the number of datasets for each resolution which greatly improves the model performance (Section 4.5).
@@
 This module is optional for the TTM backbone, particularly beneficial when the context length ($sl$) is short.
-In those scenarios, automatically detecting the resolution becomes a challenge for the model.
+In these scenarios, automatically detecting the resolution becomes a challenge for the model.
 Hence, by explicitly fusing the resolution information as a prefix, we can enhance the model’s ability to learn effectively across resolutions.
 3.2 Fine-tuning Workflow
-In the fine-tuning workflow, we deal with data from the target domain that has no overlap with the pretraining datasets.
-We have three options here: (a) In Zero-shot forecasting, we directly use the pretrained model to evaluate on the test part of the target data, (b) In Few-shot forecasting, we utilize only a tiny portion (5-10%) of the train part of the target data to quickly update the pretrained weights of the decoder and head, and subsequently, evaluate it on the test part, (c) In Full-shot forecasting, we finetune the pretrained weights of the decoder and head on the entire train part of the target data, and then, evaluate on the test part.
+In the fine-tuning workflow, we deal with data from the target domain that has no overlap with the pre-training datasets.
+We have three options here: (a) In Zero-shot forecasting, we directly use the pre-trained model to evaluate on the test part of the target data, (b) In Few-shot forecasting, we utilize only a tiny portion (5-10%) of the train part of the target data to quickly update the pre-trained weights of the decoder and head, and subsequently, evaluate it on the test part, (c) In Full-shot forecasting, we fine-tune the pre-trained weights of the decoder and head on the entire train part of the target data, and then, evaluate on the test part.
 The backbone is completely frozen during fine-tuning, and still operates in a channel-independent univariate fashion.
@@
 If pure multivariate modeling is needed, then the channel-mixer block in all the TSMixer components (see Figure 1(b)) in the decoder gets enabled to explicitly capture the channel correlation between the channels.
-The forecast head and reverse normalization perform similar operations as in the pretraining stage.
+The forecast head and reverse normalization perform similar operations as in the pre-training stage.
 The fine-tuning also optimizes the forecasting objective with MSE loss.
-This thoughtful multi-level design choice ensures our backbone excels in channel-independent pretraining, enabling effective temporal correlation modeling across diverse datasets.
+This thoughtful multi-level design choice ensures our backbone excels in channel-independent pre-training, enabling effective temporal correlation modeling across diverse datasets.
 Simultaneously, the decoder handles target-data-specific tasks like channel-correlation modeling and fine-tuning.
@@
 Let the forecast from the forecast head be $\hat{\bm{Y}}\in\mathbb{R}^{c\times fl}$.
-Let the channels $\bm{x}_{0},\cdots,\bm{x}_{N_{1}}$ denote the target variables and $\bm{x}_{N_{1}+1},\cdots,\bm{x}_{N_{1}+N_{2}}$ denote all exogenous variables with their future values known.
-First, we replace the forecast values for the exogenous channels with the true future values ($\bm{Y}$) and transpose it: $\hat{\bm{Y}}_{e}=\left[\hat{\bm{y}}_{0},\cdots,\hat{\bm{y}}_{N_{1}},\bm{y}_{N_{1}+1},\cdots,\bm{y}_{N_{1}+N_{2}}\right]\in\mathbb{R}^{fl\times c}$.
+Let the channels $\bm{x}_{0},\cdots,\bm{x}_{c^{\prime}}$ denote the target variables and $\bm{x}_{c^{\prime}+1},\cdots,\bm{x}_{c}$ denote all exogenous variables with their future values known.
+First, we replace the forecast values for the exogenous channels with the true future values ($\bm{Y}$) and transpose it: $\hat{\bm{Y}}_{e}=\left[\hat{\bm{y}}_{0},\cdots,\hat{\bm{y}}_{c^{\prime}},\bm{y}_{c^{\prime}+1},\cdots,\bm{y}_{c}\right]\in\mathbb{R}^{fl\times c}$.
 Next, to learn inter-channel lagged correlations, we patch $\hat{\bm{Y}}_{e}$ into a series of overlapped windows (i.e., patching with stride$=1$) to create a new tensor: $\hat{\bm{Y}}_{e,p}\in\mathbb{R}^{fl\times\Delta\times c}$, where $\Delta=2\cdot l+1$ with $l$ being the context length to incorporate on either side of a time point [^4] This needs padding $\hat{\bm{Y}}_{e}$ with zeros of length $l$ on both sides..
@@
 Thus, the lagged dependency of the forecasts for the target channels on the exogenous channels is seamlessly learned.
-Finally, we attach a linear head to produce the forecasts for the target channels which is then reshaped as $\hat{\bm{Y}}\in\mathbb{R}^{N_{1}\times fl}$.
+Finally, we attach a linear head to produce the forecasts for the target channels which is then reshaped as $\hat{\bm{Y}}\in\mathbb{R}^{c^{\prime}\times fl}$.
 Figure 1(c) depicts this procedure.
-Zero-shot Few-shot (5%) $fl$ TTM TTM GPT4TS PatchTST TSMixer TimesNet Dlinear FEDFormer Autoformer ETTH1 96 0.365 0.366 0.543 0.557 0.554 0.892 0.547 0.593 0.681 192 0.393 0.391 0.748 0.711 0.673 0.940 0.720 0.652 0.725 336 0.415 0.421 0.754 0.816 0.678 0.945 0.984 0.731 0.761 720 0.538 - - - - - - - - ETTH2 96 0.285 0.282 0.376 0.401 0.348 0.409 0.442 0.390 0.428 192 0.341 0.338 0.418 0.452 0.419 0.483 0.617 0.457 0.496 336 0.383 0.383 0.408 0.464 0.389 0.499 1.424 0.477 0.486 720 0.441 - - - - - - - - ETTM1 96 0.413 0.359 0.386 0.399 0.361 0.606 0.332 0.628 0.726 192 0.476 0.402 0.440 0.441 0.411 0.681 0.358 0.666 0.750 336 0.553 0.424 0.485 0.499 0.467 0.786 0.402 0.807 0.851 720 0.737 0.575 0.577 0.767 0.677 0.796 0.511 0.822 0.857 ETTM2 96 0.187 0.174 0.199 0.206 0.200 0.220 0.236 0.229 0.232 192 0.261 0.240 0.256 0.264 0.265 0.311 0.306 0.394 0.291 336 0.323 0.299 0.318 0.334 0.314 0.338 0.380 0.378 0.478 720 0.436 0.407 0.46 0.454 0.410 0.509 0.674 0.523 0.553 Weather 96 0.154 0.152 0.175 0.171 0.188 0.207 0.184 0.229 0.227 192 0.203 0.198 0.227 0.230 0.234 0.272 0.228 0.265 0.278 336 0.256 0.250 0.286 0.294 0.287 0.313 0.279 0.353 0.351 720 0.329 0.326 0.366 0.384 0.365 0.400 0.364 0.391 0.387 Electricity 96 0.169 0.142 0.143 0.145 0.147 0.315 0.150 0.235 0.297 192 0.196 0.162 0.159 0.163 0.172 0.318 0.163 0.247 0.308 336 0.209 0.184 0.179 0.175 0.190 0.340 0.175 0.267 0.354 720 0.264 0.228 0.233 0.219 0.280 0.635 0.219 0.318 0.426 Traffic 96 0.518 0.401 0.419 0.404 0.408 0.854 0.427 0.670 0.795 192 0.548 0.425 0.434 0.412 0.421 0.894 0.447 0.653 0.837 336 0.55 0.437 0.449 0.439 0.477 0.853 0.478 0.707 0.867 720 0.605 - - - - - - - - TTM Zero-shot % IMP 1% 3% 2% 31% 6% 25% 32% TTM 5% Few-shot % IMP 12% 14% 12% 38% 17% 32% 38% Table 1: Zero-shot and Few-shot 5% performance of TTM and all SOTA models on D1 datasets for varying forecast lengths ($fl$).
+Zero-shot Few-shot (5%) $fl$ TTM TTM GPT4TS PatchTST TSMixer TimesNet Dlinear FEDFormer Autoformer ETTH1 96 0.365 0.366 0.543 0.557 0.554 0.892 0.547 0.593 0.681 192 0.393 0.391 0.748 0.711 0.673 0.940 0.720 0.652 0.725 336 0.415 0.421 0.754 0.816 0.678 0.945 0.984 0.731 0.761 720 0.538 - - - - - - - - ETTH2 96 0.285 0.282 0.376 0.401 0.348 0.409 0.442 0.390 0.428 192 0.341 0.338 0.418 0.452 0.419 0.483 0.617 0.457 0.496 336 0.383 0.383 0.408 0.464 0.389 0.499 1.424 0.477 0.486 720 0.441 - - - - - - - - ETTM1 96 0.413 0.359 0.386 0.399 0.361 0.606 0.332 0.628 0.726 192 0.476 0.402 0.440 0.441 0.411 0.681 0.358 0.666 0.750 336 0.553 0.424 0.485 0.499 0.467 0.786 0.402 0.807 0.851 720 0.737 0.575 0.577 0.767 0.677 0.796 0.511 0.822 0.857 ETTM2 96 0.187 0.174 0.199 0.206 0.200 0.220 0.236 0.229 0.232 192 0.261 0.240 0.256 0.264 0.265 0.311 0.306 0.394 0.291 336 0.323 0.299 0.318 0.334 0.314 0.338 0.380 0.378 0.478 720 0.436 0.407 0.46 0.454 0.410 0.509 0.674 0.523 0.553 Weather 96 0.154 0.152 0.175 0.171 0.188 0.207 0.184 0.229 0.227 192 0.203 0.198 0.227 0.230 0.234 0.272 0.228 0.265 0.278 336 0.256 0.250 0.286 0.294 0.287 0.313 0.279 0.353 0.351 720 0.329 0.326 0.366 0.384 0.365 0.400 0.364 0.391 0.387 Electricity 96 0.169 0.142 0.143 0.145 0.147 0.315 0.150 0.235 0.297 192 0.196 0.162 0.159 0.163 0.172 0.318 0.163 0.247 0.308 336 0.209 0.184 0.179 0.175 0.190 0.340 0.175 0.267 0.354 720 0.264 0.228 0.233 0.219 0.280 0.635 0.219 0.318 0.426 Traffic 96 0.518 0.401 0.419 0.404 0.408 0.854 0.427 0.670 0.795 192 0.548 0.425 0.434 0.412 0.421 0.894 0.447 0.653 0.837 336 0.55 0.437 0.449 0.439 0.477 0.853 0.478 0.707 0.867 720 0.605 - - - - - - - - TTM Zero-shot % IMP 1% 3% 2% 31% 6% 25% 32% TTM 5% Few-shot % IMP 12% 14% 12% 38% 17% 32% 38% Table 1: Zero-shot and Few-shot 5% performance (MSE) of TTM and all SOTA models on D1 datasets for varying forecast lengths ($fl$).
 Bold and underscore denote the best and the second best.
@@
 Metric Data GPT4TS TTM (Ours) IMP (X) FT NPARAMS (M) Weather 3.2 0.29 (14X) Electricity 3.2 0.29 Traffic 5.5 0.29 TL NPARAMS (M) Weather 84.3 0.8 (106X) Electricity 84.3 0.8 Traffic 86.6 0.8 EPOCH TIME (s) Weather 37 1 (65X) Electricity 241 4 Traffic 782 8 MAX MEMORY (GB) Weather 21.9 1.06 (27X) Electricity 39.7 2.02 Traffic 53.7 1.34 TEST TIME (s) Weather 52.53 1.66 (54X) Electricity 353.4 8.07 Traffic 1163.9 13.5 Table 2: Computational Improvement in few-shot 10% for $fl$=96 using 1 A100 GPU.
-nX denotes n times average improvement across datasets (Avg.
-IMP).
+nX denotes n times average improvement across datasets (IMP).
 10% 25% 50% 75% 100% IMP TTM 0.422 0.421 0.413 0.402 0.398 - SimMTM 0.591 0.535 0.491 0.466 0.415 17% Ti-MAE 0.660 0.594 0.55 0.475 0.466 24% TST 0.783 0.641 0.525 0.516 0.469 28% LaST 0.645 0.610 0.540 0.479 0.443 23% TF-C 0.799 0.736 0.731 0.697 0.635 43% CoST 0.784 0.624 0.540 0.494 0.428 26% TS2Vec 0.655 0.632 0.599 0.577 0.517 31%
-Figure 2: Cross transfer learning improvement (IMP) for self-supervised pretraining methods in various few-shot settings (10%,25%,50%,75%,100%).
+Figure 2: Cross transfer learning MSE improvement (IMP) for self-supervised pre-training methods in various few-shot settings (10%,25%,50%,75%,100%).
 4 Experiments and Results
 4.1 Experimental Setting
-Datasets: Pretraining employs a subset of the Monash data hub Godahewa et al.
-(2021) of size $\sim 240$M samples.
-We specifically exclude datasets that consist of yearly, monthly, and daily series as they do not possess sufficient length for the long-term forecasting task.
-Moreover, we remove all the datasets that we utilize for evaluation.
-This includes weather, electricity, and traffic datasets.
-For zero/few/full-shot evaluation we consider seven datasets (D1): ETTH1, ETTH2, ETTM1, ETTM2, Weather, Electricity (ECL), Traffic as popularly used in most prior SOTA works Zhou et al.
+Datasets: Pre-training employs a subset of the Monash data hub Godahewa et al.
+(2021) of size $\sim 244$M samples.
+We specifically exclude datasets (like yearly, monthly) as they do not possess sufficient length for the long-term forecasting task.
+Moreover, we remove all the datasets that we utilize for evaluation (i.e., weather, electricity, and traffic).
+For zero/few-shot evaluation we consider seven public datasets (D1): ETTH1, ETTH2, ETTM1, ETTM2, Weather, Electricity & Traffic as popularly used in most prior SOTA works Zhou et al.
 (2021); Nie et al.
@@
 Since these datasets do not contain any exogenous variables nor exhibit cross-channel correlation benefits, we incorporate four other datasets (D2) for separately validating the efficacy of the decoder channel mixing and exogenous mixer module: bike sharing (BS) Fanaee-T (2013), carbon capture plant (CC) Jablonka et al.
-(2023), and two more datasets from Business-IT observability domain Biz (2023): Application (APP) and Service (SER).
-Details are provided in the Appendix.
-SOTA Bechmarks: We benchmark TTM model with the latest publicly available SOTA models for multivariate forecasting categorized as follows: [label=()] 1.
-LLM-based TS pretrained models:GPT4TS Zhou et al.
+(2023), and 2 more datasets from Biz-IT observability domain ITB (2023): Application (APP) and Service (SER).
+Refer Appendix for full data details.
+SOTA Bechmarks: We benchmark TTM with the latest public SOTA forecasting models categorized as follows: [label=()] 1.
+LLM-based TS pre-trained models:GPT4TS Zhou et al.
 (2023), LLMTime Gruver et al.
 (2023) 2.
-Self-supervised pretrained models: SimMTM Dong et al.
+Self-supervised pre-trained models: SimMTM Dong et al.
 (2023),Ti-MAE Li et al.
@@
 (2022)
-Zero-shot Few-shot (10%) Data TTM (Ours) TTM (Ours) GPT4TS PatchTST TSMixer TimesNet Dlinear FEDFormer Autoformer ETTH1 0.428 0.422 0.590 0.633 0.578 0.869 0.691 0.639 0.702 ETTH2 0.362 0.360 0.397 0.415 0.370 0.479 0.605 0.466 0.488 ETTM1 0.545 0.458 0.464 0.501 0.441 0.677 0.411 0.722 0.802 ETTM2 0.302 0.278 0.293 0.296 0.282 0.320 0.316 0.463 1.342 Electricity 0.210 0.180 0.176 0.176 0.180 0.323 0.180 0.346 0.431 Traffic 0.555 0.428 0.440 0.430 0.429 0.951 0.446 0.663 0.749 Weather 0.236 0.227 0.238 0.242 0.236 0.279 0.241 0.284 0.300 TTM Zero-shot improvements -4% -2% -7% 27% 2% 27% 39% TTM Few-shot (10%) improvements 7% 9% 4% 34% 13% 34% 45% Table 3: Zero-shot and Few-shot 10% performance of TTM and all SOTA models on seven datasets averaged across 4 forecast lengths.
-The numbers with bold and underscore denote the best the second best results respectively.
+Zero-shot Few-shot (10%) Data TTM (Ours) TTM (Ours) GPT4TS PatchTST TSMixer TimesNet Dlinear FEDFormer Autoformer ETTH1 0.428 0.422 0.590 0.633 0.578 0.869 0.691 0.639 0.702 ETTH2 0.362 0.360 0.397 0.415 0.370 0.479 0.605 0.466 0.488 ETTM1 0.545 0.458 0.464 0.501 0.441 0.677 0.411 0.722 0.802 ETTM2 0.302 0.278 0.293 0.296 0.282 0.320 0.316 0.463 1.342 Electricity 0.210 0.180 0.176 0.176 0.180 0.323 0.180 0.346 0.431 Traffic 0.555 0.428 0.440 0.430 0.429 0.951 0.446 0.663 0.749 Weather 0.236 0.227 0.238 0.242 0.236 0.279 0.241 0.284 0.300 TTM Zero-shot improvements -4% -2% -7% 27% 2% 27% 39% TTM Few-shot (10%) improvements 7% 9% 4% 34% 13% 34% 45% Table 3: Zero-shot and Few-shot 10% performance (MSE) of TTM over all SOTA models on D1 datasets, averaged across 4 forecast lengths.
+The numbers with bold and underscore denote the best and the second best results respectively.
 The full table is in Appendix.
-Data $fl$ Zero-shot LLMTime TTM Traffic 96 0.340 0.195 192 0.526 0.367 Weather 96 0.107 0.031 192 0.062 0.067 Electricity 96 0.609 0.238 192 0.960 0.488 ETTM2 96 0.167 0.164 192 0.198 0.228 TTM IMP (%) 29% Table 4: Comparison between TTM and LLMTime in zero-shot setting.
-TTM Model Details: By default, the following data and model configuration is used for pretraining: input sequence length $sl$ = 512, patch length $pl$ = 64, stride $s$ = 64, number of patches $n=512/64=8$, forecast length $fl\in\{96,192,336,720\}$, number of levels $L$ = 6, number of TTM blocks per level $M$ = 2, batch size $b$ = 3000, number of epochs $ep$ = 25, and dropout $do$ = 0.2.
-TSMixer specific hyperparameters include feature scaler $fs$ = 3, hidden feature size $hf$ = $fs*pl=192$, expansion feature size $ef=hf*2=384$.
-Please note that, $hf$ and $n$ will change across TTM blocks based on the adaptive patching strategy, and the decoder works in channel-independent mode during pretraining.
-Resolution prefix tuning is disabled by default and enabled only for shorter context length (as explained in Table.
+Data $fl$ Zero-shot LLMTime TTM Traffic 96 0.340 0.195 192 0.526 0.367 Weather 96 0.107 0.031 192 0.062 0.067 Electricity 96 0.609 0.238 192 0.960 0.488 ETTM2 96 0.167 0.164 192 0.198 0.228 TTM IMP (%) 29% Table 4: TTM vs LLMTime MSE Improvement (IMP) in zero-shot setting.
+TTM Model Details: For our experiments, we build 5 pre-trained models for the following $sl$,$fl$ configuration: (512,96), (512,192), (512, 336), (512,720) and (96,24).
+Based on the $sl$ and $fl$ requirement of the target dataset, a suitable pre-trained model is selected to initialize the weights accordingly for fine-tuning.
+Since our pre-trained models are very small, they can be trained quickly in a few hours (4-8 hrs), as opposed to several days/weeks in standard approaches.
+Hence, pre-training multiple TTM models is no longer a practical constraint.
+Pre-training is performed in a distributed fashion with 50 CPUs and 6 A100 GPUs with 40 GB GPU memory while fine-tuning uses only 1 GPU.
+Other model configurations are as follows: patch length $pl$ = 64 (for $sl$ 512) and 8 (for $sl$ 96), stride $s$ = $pl$, number of levels $L$ = 6, number of TTM blocks per level $M$ = 2, number of decoder layers = 2, batch size $b$ = 3K, number of epochs $ep$ = 20, and dropout $do$ = 0.2.
+TSMixer specific hyperparameters include feature scaler $fs$ = 3, hidden feature size $hf$ = $fs*pl$, expansion feature size $ef=hf*2$.
+Resolution prefix tuning is disabled by default and enabled only for shorter context lengths (as explained in Table.
 9).
-Training is performed in a distributed fashion with 50 CPUs and 6 NVIDIA A100 GPUs with 40 GB GPU memory.
-We build a pretrained model using Monash data for each specified $fl$.
-Since pretrained models are very small, they can be trained quickly in a few hours (4-8 hrs based on $fl$ length), as opposed to several days/weeks in standard approaches.
-For few/full-shot fine-tuning, all the above pretraining parameters remain the same except the following.
-Based on the target data characteristics, the decoder channel-mixing and/or exogenous mixer block is enabled/disabled (both enabled for D2 datasets).
-In addition, head dropout changes to 0.7 for smaller ETT datasets and 0.2 otherwise and batch size ranges between 8-64.
-All the hyperparameters are adjusted based on the validation performance, and the final test results are reported in the paper.
-Data BS CC APP SER % IMP
+Decoder channel-mixing and exogenous mixer blocks are disabled during pre-training and enabled for D2 datasets during fine-tuning.
+All other model parameters remain the same across pre-training and fine-tuning except for a few parameters like dropout and batch size that can be adjusted based on the target data.
+Hyperparameters are selected based on the validation performance, and the test results are reported.
+For full implementation details, refer to the Appendix.
+MSE is used as the default evaluation metric.
+BS CC APP SER % IMP
 TTM-CM 0.582 0.250 0.042 0.114
@@
 GPT4TS 0.645 0.254 0.075 0.135 18%
-Table 5: Effect of Decoder Mixing and Exog Fusion. Results are reported using $(sl,fl)$ with values of $(512,96)$ for BS and $(96,24)$ for other datasets.
+Table 5: Effect of Decoder Mixing and Exog Fusion. MSE Results are reported using $(sl,fl)$ with values of $(512,96)$ for BS dataset and $(96,24)$ for other D2 datasets.
 4.2 TTM’s Zero/Few-shot Performance
-Table 1 compares the performance of TTM model in zero-shot and few-shot (5%) settings.
+Table 1 compares the performance of TTM model in zero-shot and few-shot (5%) settings across multiple $fl$s.
 Baseline results are reported from  Zhou et al.
 (2023) as we use the same few-shot data filtering strategy as followed in that paper.
-Note that for zero-shot performance, the TTM model pretrained on the Monash data is directly evaluated on the test set for every dataset.
-We report the test MSE for multiple forecast horizons varying from 96 to 720.
+Note that for zero-shot performance, the pre-trained TTM model is directly evaluated on the test set.
 The 5% few-shot TTM outperforms the SOTAs in most of the cases with significant accuracy gains (12-38%).
-An even more impressive observation is that the TTM in a zero-shot setting is also able to outperform most of the SOTAs which are trained on 5% of the target data.
-This observation establishes the generalization ability of the pretrained TTM model on the target datasets.
+An even more impressive observation is that the TTM in zero-shot setting is also able to outperform most of the SOTAs which are trained on 5% of the target data.
+This observation establishes the generalization ability of the pre-trained TTM model on the target datasets.
 Likewise, Table 3 shows the 10% few-shot performance of the TTM, where we outperform all the existing SOTAs by 4-45% accuracy gains.
@@
 In this evaluation, we outperform LLMTime by a substantial margin of 29%.
-In addition, there are alternative pretraining approaches like masked modeling and contrastive learning techniques that may not offer universal transfer learning capabilities across all datasets (like TTM), but they excel in enabling cross-transfer learning between two datasets when carefully selected.
+In addition, there are alternative pre-training approaches like masked modeling and contrastive learning techniques that may not offer universal transfer learning capabilities across all datasets (like TTM), but they excel in enabling cross-transfer learning between two datasets when carefully selected.
 Table 2 illustrates the cross-transfer learning from ETTH2 to ETTH1 for these models in various few-shot settings (as reported in  Dong et al.
@@
 Thus, TTM significantly beats all the existing SOTAs including the recent popular transfer learning benchmarks based on LLMs.
-The notable part here is that we get these substantial accuracy gains with significantly reduced compute/runtime requirements as explained next.
+Notably, we achieve these accuracy gains with significantly reduced compute requirements, as elaborated next.
 Data Zero-shot Few-shot (5%)) Few-shot (10%))
@@
 IMP 36% 21% 12% 9% 2%
-Table 6: Effect of pretraining (PT) over Randomly Initialized (RI) weights, averaged across 4 values of $fl$. Full table in Appendix.
-Data Zero-shot Few-shot (10%) Vanilla Adaptive Patching Vanilla Adaptive Patching ETTH1 0.501 0.428 0.478 0.422 ETTH2 0.375 0.362 0.372 0.360 ETTM1 0.577 0.545 0.463 0.458 ETTM2 0.310 0.302 0.272 0.278 Electricity 0.215 0.210 0.178 0.180 Traffic 0.556 0.555 0.431 0.428 Weather 0.239 0.236 0.230 0.227 IMP (%) 4% 2% Table 7: Effect of adaptive Patching, averaged across 4 values of $fl$.
-Data Zero-shot Original Data With Downsample ETTH1 0.603 0.379 ETTH2 0.351 0.313 ETTM1 0.852 0.444 ETTM2 0.261 0.224 Electricity 0.424 0.182 Traffic 0.898 0.533 Weather 0.186 0.178 IMP (%) 30% Table 8: Effect of downsampling, averaged across $fl$ 96 and 192.
-Data $HL$ = 96 $HL$ = 512 NP P NP P ETTH1 0.373 0.358 0.365 0.360 ETTH2 0.180 0.179 0.285 0.280 ETTM1 0.559 0.387 0.413 0.384 ETTM2 0.127 0.108 0.187 0.194 Weather 0.103 0.103 0.154 0.160 Electricity 0.208 0.201 0.169 0.175 Traffic 0.754 0.740 0.518 0.518 IMP (%) 8% 0% Table 9: Resolution Prefix Tuning: With (P) vs Without(NP)
+Table 6: Improvement (IMP) of Pre-Training (PT) over Randomly Initialized (RI) weights. MSE averaged across all $fl$s.
+Data Zero-shot Few-shot (10%) Vanilla Adaptive Patching Vanilla Adaptive Patching ETTH1 0.501 0.428 0.478 0.422 ETTH2 0.375 0.362 0.372 0.360 ETTM1 0.577 0.545 0.463 0.458 ETTM2 0.310 0.302 0.272 0.278 Electricity 0.215 0.210 0.178 0.180 Traffic 0.556 0.555 0.431 0.428 Weather 0.239 0.236 0.230 0.227 IMP (%) 4% 2% Table 7: Effect of adaptive Patching, MSE averaged across 4 values of $fl$
+Data Zero-shot Original Data With Downsample ETTH1 0.603 0.379 ETTH2 0.351 0.313 ETTM1 0.852 0.444 ETTM2 0.261 0.224 Electricity 0.424 0.182 Traffic 0.898 0.533 Weather 0.186 0.178 IMP (%) 30% Table 8: Downsampling effect, MSE averaged across $fl$s: 96, 192
+Data $HL$ = 96 $HL$ = 512 NP P NP P ETTH1 0.373 0.358 0.365 0.360 ETTH2 0.180 0.179 0.285 0.280 ETTM1 0.559 0.387 0.413 0.384 ETTM2 0.127 0.108 0.187 0.194 Weather 0.103 0.103 0.154 0.160 Electricity 0.208 0.201 0.169 0.175 Traffic 0.754 0.740 0.518 0.518 IMP (%) 8% 0% Table 9: Resolution Prefix Tuning:With (P) vs Without(NP).
+MSE reported
 4.3 Computational Benefits of TTM
-Table  2 compares the computational benefits of TTM with GPT4TS, the most popular LLM based TS model and the current best SOTA in few-shot setting.
-We run both models in a single A100 GPU with matching batch sizes for a fair comparison.
-Multi-GPU usage is deliberately avoided in this experiment to eliminate IPC overheads and ensure precise metric calculations.
-Since GPT4TS is based on GPT-2, it consumes a huge model footprint and exhibits slow execution as compared to TTM.
-In specific, TTM achieves 14X cut in Finetune (FT) parameters (NPARAMS) and a remarkable 106X reduction in Total (TL) parameters.
+Table  2 compares the computational benefits of TTM over GPT4TS, the most popular LLM-TS model and the current best SOTA in few-shot setting.
+For a fair comparison, we run both models with the best-reported parameters in a single A100 GPU with the same setup (Details in Appendix).
+Since GPT4TS is based on GPT-2, it consumes a huge model footprint and exhibits slow execution.
+In specific, TTM achieves 14X cut in Fine-tune (FT) parameters (NPARAMS) and a remarkable 106X reduction in Total (TL) parameters.
 This reduction in model footprint further leads to a significant reduction in fine-tuning EPOCH TIME by 65X, MAX MEMORY usage by 27X, and total inference time on the entire test data (TEST TIME) by 54X.
-The computational advantages also apply over other LLM-based time series models like LLMTime, built on LLaMA, which is larger than GPT-2 (utilized in GPT4TS).
-In fact, LLMTime used a very small test set on these datasets as opposed to the standard test set to overcome this slow-execution behavior.
-4.4 TTM’s effectiveness in cross-channel and exogenous/control modeling
+The computational benefits also apply over other LLM-based time series models like LLMTime, built on LLaMA, which is larger than GPT-2.
+In fact, LLMTime used a very small test set on these datasets as opposed to the standard test set to overcome this slow-execution.
+4.4 TTM’s effectiveness in cross-channel and exogenous modeling
 Since the datasets ($D1$) used in previous experiments do not have exogenous variables, we evaluate the effectiveness of TTM on 4 other datasets ($D2$, as explained in  4.1) to quantify its benefits.
-Table 5 shows the performance of the Monash-pretrained TTM model fine-tuned on the target data with exogenous mixer module and decoder channel-mixing enabled (TTM-CM).
-We compare the full-shot performance of TTM-CM with the pretrained TTM used in zero-shot setting (TTM-Zero-shot), and plain TTM (TTM) and other primary SOTAs (PatchTST, TSMixer variants, and GPT4TS) trained from scratch.
-Remaining SOTAs are not reported in Table 5 considering their inferior performance and space constraints.
+Since these datasets are already very small, we used their full data for fine-tuning.
+Table 5 shows the performance of the pre-trained TTM model fine-tuned on the target data with exogenous mixer module and decoder channel-mixing enabled (TTM-CM).
+We compare the performance of TTM-CM with the pre-trained TTM used in zero-shot setting (TTM-Zero-shot), and plain TTM (TTM) and other primary SOTAs (PatchTST, TSMixer variants, and GPT4TS) trained from scratch.
+Other SOTAs are not reported here, considering their inferior performance and space constraints.
 Specifically, we compare with TSMixer with channel-mixing enabled (TSMixer-CM) and TSMixer with Cross-Channel reconciliation head (TSMixer-CC) Ekambaram et al.
-(2023) as they are latest SOTAs in channel-correlation modeling.
-We can see that TTM-CM outperforms all the competitive models with a significant margin (15-44%), thus, demonstrating that the power of TTM in capturing inter-channel correlations that are prevalent in purely multivariate datasets.
+(2023) as they are the latest SOTAs in channel-correlation modeling.
+We can see that TTM-CM outperforms all the competitive models with a significant margin (15-44%), thus, demonstrating the power of TTM in capturing inter-channel correlations.
 4.5 Ablation Studies
-Effect of Pretraining: Table 6 illustrates the advantages of the proposed pretraining (PT) approach in comparison to randomly initialized (RI) model weights.
-In the case of zero-shot, the pretrained TTM (PT) exhibits 36% improvement over RI.
+Effect of Pre-training: Table 6 illustrates the advantages of the proposed pre-training (PT) approach in comparison to randomly initialized (RI) model weights.
+In the case of zero-shot, the pre-trained TTM (PT) exhibits 36% improvement over RI.
 This outcome is expected, as random weights are directly employed for forecasting with RI.
 For the 5% few-shot scenario, PT achieves noteworthy improvements of 21% and 12% over RI when fine-tuned for 5 and 50 epochs, respectively.
-This underscores the utility of pretrained weights in facilitating quick learning with a limited amount of target data.
+This underscores the utility of pre-trained weights in facilitating quick learning with a limited amount of target data.
 Even in the case of 10% few-shot, PT continues to demonstrate an improved performance of 9% for 5 epochs and 2% for 50 epochs.
-Thus, the performance impact of leveraging pretrained weights greatly increases as the size of the training data and the available time-to-finetune reduces.
+Thus, the performance impact of leveraging pre-trained weights greatly increases as the size of the training data and the available time-to-fine-tune reduces.
 Effect of Adaptive Patching: The Adaptive patching presented in Section 3.1.1 is a key component of the TTM backbone.
-Table 7 demonstrates the relative improvements in MSE when we employ adaptive patching with respect to a vanilla TTM backbone without this module.
+Table 7 demonstrates the relative improvements when we employ adaptive patching with respect to a vanilla TTM backbone without this module.
 We can see that adaptive patching adds an extra 4% improvement during zero-shot and 2% improvement for 10% few-shot.
-MSE improvements are observed in almost all datasets which justifies the effectiveness of the adaptive patching module when evaluated on multiple datasets of different resolutions.
+Performance improvements are observed in almost all datasets which justifies the effectiveness of the adaptive patching module when evaluated on multiple datasets of different resolutions.
 Effect of Augmentation via Downsampling: Table 8 compares the zero-shot test performances between the TTM model trained on the original Monash data and the TTM trained on the augmented Monash data after downsampling as described in Section 3.1.
-It is evident that there is a significant improvement of 30% for the model pretrained on the augmented datasets as it adds more data coverage across various resolutions.
+There is a significant improvement of 30% for the model pre-trained on the augmented datasets as it adds more data coverage across various resolutions.
 This observation highlights the strength of the proposed TTM model to learn from diverse datasets together, and increasing the number of datasets for each resolution helps.
@@
 5 Conclusions and Future Work
-Considering the diversity and limited availability of public TS pretraining data, pretraining large models for effective transfer learning poses several challenges in time series.
-Hence, we propose TTM, a multi-level Tiny Time Mixer model designed for efficient pretraining on limited diverse multi-resolution datasets.
+Considering the diversity and limited availability of public TS pre-training data, pre-training large models for effective transfer learning poses several challenges in time series.
+Hence, we propose TTM, a multi-level Tiny Time Mixer model designed for efficient pre-training on limited diverse multi-resolution datasets.
 TTM achieves state-of-the-art results in zero/few-shot forecasting, offering significant computational efficiency and supporting cross-channel and exogenous variables — critical features lacking in existing popular methods.

### 09 Apr 2024 v5

@@
 Multivariate time series (TS) forecasting entails predicting future values for multiple interrelated time series based on their historical data.
-This field has advanced significantly, applying statistical and machine learning (ML) methods Hyndman and Athanasopoulos (2021) across domains like weather, traffic, retail, and energy.
+This field has advanced significantly, applying statistical and machine learning (ML) methods [?] across domains like weather, traffic, retail, and energy.
 In general, each time series represents a variable or channel111“Channel” refers to the individual time series in multivariate data (i.e., a multivariate TS is a multi-channel signal)..
@@
 We term these non-forecasting variables as exogenous, and the variables requiring forecast as target variables.
-Related Work: Recent advances in multivariate forecasting have been marked by the advent of transformer-based Vaswani et al.
-(2017) approaches, exemplified by models like PatchTST Nie et al.
-(2023), Autoformer Wu et al.
-(2021), Informer Zhou et al.
-(2021), and FEDFormer Zhou et al.
-(2022).
+Related Work: Recent advances in multivariate forecasting have been marked by the advent of transformer-based [?] approaches, exemplified by models like PatchTST [?], Autoformer [?], Informer [?], and FEDFormer [?].
 These models have demonstrated notable improvements over traditional statistical and ML methods.
-Furthermore, architectures based on MLP-Mixer Tolstikhin et al.
-(2021), such as TSMixer Ekambaram et al.
-(2023), have emerged as efficient transformer alternatives, boasting 2-3X reduced compute and memory requirements with no accuracy compromise compared to their transformer counterparts.
+Furthermore, architectures based on MLP-Mixer [?], such as TSMixer [?], have emerged as efficient transformer alternatives, boasting 2-3X reduced compute and memory requirements with no accuracy compromise compared to their transformer counterparts.
 However, none of these advanced approaches have successfully demonstrated the ability to create general pre-trained models that can successfully transfer the learning to unseen target TS dataset, in a similar way as popularly witnessed in NLP and vision tasks.
 This is very challenging in the TS domain due to the diverse nature of the datasets across applications and the limited public availability of TS data for pre-training.
-There are existing self-supervised pre-training TS approaches using masked modeling and contrastive learning techniques such as SimMTM Dong et al.
-(2023) and TF-C Zhang et al.
-(2022) that offer transfer learning between two datasets when carefully selected based on the dataset properties.
+There are existing self-supervised pre-training TS approaches using masked modeling and contrastive learning techniques such as SimMTM [?] and TF-C [?] that offer transfer learning between two datasets when carefully selected based on the dataset properties.
 However, they fail to provide universal transfer learning capabilities across datasets.
 Consequently, there has been a recent growing trend to employ pre-trained large language models (LLMs) for TS forecasting, treating it as a cross-domain transfer learning task.
-These universal cross-transfer approaches, specifically recent works such as LLMTime Gruver et al.
-(2023) and GPT4TS Zhou et al.
-(2023) yield promising results in few/zero-shot forecasting approaches.
+These universal cross-transfer approaches, specifically recent works such as LLMTime [?] and GPT4TS [?] yield promising results in few/zero-shot forecasting approaches.
 These models are bootstrapped from GPT-2/3 or LLAMA-2 with suitable tokenization strategies to adapt to time-series domains.
@@
 Toward this, we propose Multi-level Tiny Time Mixers (TTM), a significantly smaller model ($\leq$1M parameters) based on the lightweight TSMixer architecture, exclusively trained on diverse TS corpora for effective zero/few-shot multivariate TS forecasting via transfer learning.
-In particular, TTM is pre-trained using multiple public datasets ($\sim$ 244M samples) from the Monash data repository [^2] Accessible at https://forecastingdata.org/ Godahewa et al.
-(2021)).
+In particular, TTM is pre-trained using multiple public datasets ($\sim$ 244M samples) from the Monash data repository [^2] Accessible at https://forecastingdata.org/ [?]).
 Note that the datasets exhibit considerable diversity in terms of characteristics, such as the different domains, temporal resolution [^3] Resolution refers to the sampling rate of the input time series (e.g., hourly, 10 minutes, 15 minutes, etc.) (spanning from second to daily), lengths, and number of channels.
@@
 2.1 Multi-level Modeling
-TTM follows a multi-level architecture consisting of four key components (see Figure 1(a)): (1) The TTM Backbone is assembled using building blocks derived from the efficient TSMixer architecture Ekambaram et al.
-(2023).
+TTM follows a multi-level architecture consisting of four key components (see Figure 1(a)): (1) The TTM Backbone is assembled using building blocks derived from the efficient TSMixer architecture [?].
 TSMixer is based on simple MLP blocks that enable mixing of features within patches, across patches and channels, surpassing existing transformer-based TS approaches with minimal computational requirements.
@@
 2.2 Pre-processing
-As shown in Figure 1(a) with colorless blocks, the historical time series $\bm{X}$ is first normalized per instance to have zero mean and unit standard deviation for each channel dimension, to tackle any possible distribution shifts Nie et al.
-(2023); Ekambaram et al.
-(2023).
+As shown in Figure 1(a) with colorless blocks, the historical time series $\bm{X}$ is first normalized per instance to have zero mean and unit standard deviation for each channel dimension, to tackle any possible distribution shifts [??].
 This process is reversed at the end before computing the loss.
 The normalized data $\overline{\bm{X}}$ is subsequently patched $\bm{X}_{p}\in\mathbb{R}^{c\times n\times pl}$ into $n$ non-overlapping windows, each of length $pl$ and then, passed to the TTM backbone.
-Patching, as introduced in Nie et al.
-(2023), has proven to be highly valuable for forecasting.
+Patching, as introduced in [?], has proven to be highly valuable for forecasting.
 Its effectiveness lies in preserving local semantic information, accommodating longer history, and reducing computation.
@@
 3.1 Pre-training Workflow
-In the pre-training stage, we train the model on a large collection of public datasets from the Monash data repositoryGodahewa et al.
-(2021).
+In the pre-training stage, we train the model on a large collection of public datasets from the Monash data repository[?].
 Since the primary focus of TTM is forecasting, pre-training is modeled with a direct forecasting objective.
@@
 Thus for a given input context length $sl$ and forecast length $fl$, we get a pre-trained model capturing the common temporal forecasting dynamics and seasonal patterns as observed in the overall pre-training data.
-3.1.1 Multi-Resolution Pre-training via TTM Backbone
+Multi-Resolution Pre-training via TTM Backbone
 The majority of the pre-training happens in the TTM backbone.
@@
 This idea of adaptive patching is popular and very successful in the vision domain (E.g.
-Swin transformers Liu et al.
-(2021) and we are the first to port it successfully to the time-series domain to resolve multi-resolution issues in pre-training with diverse TS datasets.
+Swin transformers [?] and we are the first to port it successfully to the time-series domain to resolve multi-resolution issues in pre-training with diverse TS datasets.
 Figure 1(b) shows the TTM backbone for $L=3$ and $M=2$.
@@
 Resolution Prefix Tuning: This technique explicitly learns and incorporates a new patch embedding as a prefix into the input data based on the input resolution type (see Figure 1(b)).
-Similar to the concept of prefix tuning Li and Liang (2021), this approach provides an explicit signal to the model about the resolution type for resolution-conditioned modeling.
+Similar to the concept of prefix tuning [?], this approach provides an explicit signal to the model about the resolution type for resolution-conditioned modeling.
 First, we map every resolution to a unique integer, which is then passed through an embedding layer to project it to the hidden dimension, $hf$.
@@
 Figure 1(c) depicts this procedure.
-Zero-shot Few-shot (5%) $fl$ TTM TTM GPT4TS PatchTST TSMixer TimesNet Dlinear FEDFormer Autoformer ETTH1 96 0.365 0.366 0.543 0.557 0.554 0.892 0.547 0.593 0.681 192 0.393 0.391 0.748 0.711 0.673 0.940 0.720 0.652 0.725 336 0.415 0.421 0.754 0.816 0.678 0.945 0.984 0.731 0.761 720 0.538 - - - - - - - - ETTH2 96 0.285 0.282 0.376 0.401 0.348 0.409 0.442 0.390 0.428 192 0.341 0.338 0.418 0.452 0.419 0.483 0.617 0.457 0.496 336 0.383 0.383 0.408 0.464 0.389 0.499 1.424 0.477 0.486 720 0.441 - - - - - - - - ETTM1 96 0.413 0.359 0.386 0.399 0.361 0.606 0.332 0.628 0.726 192 0.476 0.402 0.440 0.441 0.411 0.681 0.358 0.666 0.750 336 0.553 0.424 0.485 0.499 0.467 0.786 0.402 0.807 0.851 720 0.737 0.575 0.577 0.767 0.677 0.796 0.511 0.822 0.857 ETTM2 96 0.187 0.174 0.199 0.206 0.200 0.220 0.236 0.229 0.232 192 0.261 0.240 0.256 0.264 0.265 0.311 0.306 0.394 0.291 336 0.323 0.299 0.318 0.334 0.314 0.338 0.380 0.378 0.478 720 0.436 0.407 0.46 0.454 0.410 0.509 0.674 0.523 0.553 Weather 96 0.154 0.152 0.175 0.171 0.188 0.207 0.184 0.229 0.227 192 0.203 0.198 0.227 0.230 0.234 0.272 0.228 0.265 0.278 336 0.256 0.250 0.286 0.294 0.287 0.313 0.279 0.353 0.351 720 0.329 0.326 0.366 0.384 0.365 0.400 0.364 0.391 0.387 Electricity 96 0.169 0.142 0.143 0.145 0.147 0.315 0.150 0.235 0.297 192 0.196 0.162 0.159 0.163 0.172 0.318 0.163 0.247 0.308 336 0.209 0.184 0.179 0.175 0.190 0.340 0.175 0.267 0.354 720 0.264 0.228 0.233 0.219 0.280 0.635 0.219 0.318 0.426 Traffic 96 0.518 0.401 0.419 0.404 0.408 0.854 0.427 0.670 0.795 192 0.548 0.425 0.434 0.412 0.421 0.894 0.447 0.653 0.837 336 0.55 0.437 0.449 0.439 0.477 0.853 0.478 0.707 0.867 720 0.605 - - - - - - - - TTM Zero-shot % IMP 1% 3% 2% 31% 6% 25% 32% TTM 5% Few-shot % IMP 12% 14% 12% 38% 17% 32% 38% Table 1: Zero-shot and Few-shot 5% performance (MSE) of TTM and all SOTA models on D1 datasets for varying forecast lengths ($fl$).
+Zero-shot Few-shot (5%) $fl$ TTM TTM GPT4TS PatchTST TSMixer TimesNet Dlinear FEDFormer Autoformer ETTH1 96 0.365 0.366 0.543 0.557 0.554 0.892 0.547 0.593 0.681 192 0.393 0.391 0.748 0.711 0.673 0.940 0.720 0.652 0.725 336 0.415 0.421 0.754 0.816 0.678 0.945 0.984 0.731 0.761 720 0.538 - - - - - - - - ETTH2 96 0.285 0.282 0.376 0.401 0.348 0.409 0.442 0.390 0.428 192 0.341 0.338 0.418 0.452 0.419 0.483 0.617 0.457 0.496 336 0.383 0.383 0.408 0.464 0.389 0.499 1.424 0.477 0.486 720 0.441 - - - - - - - - ETTM1 96 0.413 0.359 0.386 0.399 0.361 0.606 0.332 0.628 0.726 192 0.476 0.402 0.440 0.441 0.411 0.681 0.358 0.666 0.750 336 0.553 0.424 0.485 0.499 0.467 0.786 0.402 0.807 0.851 720 0.737 0.575 0.577 0.767 0.677 0.796 0.511 0.822 0.857 ETTM2 96 0.187 0.174 0.199 0.206 0.200 0.220 0.236 0.229 0.232 192 0.261 0.240 0.256 0.264 0.265 0.311 0.306 0.394 0.291 336 0.323 0.299 0.318 0.334 0.314 0.338 0.380 0.378 0.478 720 0.436 0.407 0.46 0.454 0.410 0.509 0.674 0.523 0.553 Weather 96 0.154 0.152 0.175 0.171 0.188 0.207 0.184 0.229 0.227 192 0.203 0.198 0.227 0.230 0.234 0.272 0.228 0.265 0.278 336 0.256 0.250 0.286 0.294 0.287 0.313 0.279 0.353 0.351 720 0.329 0.326 0.366 0.384 0.365 0.400 0.364 0.391 0.387 Electricity 96 0.169 0.142 0.143 0.145 0.147 0.315 0.150 0.235 0.297 192 0.196 0.162 0.159 0.163 0.172 0.318 0.163 0.247 0.308 336 0.209 0.184 0.179 0.175 0.190 0.340 0.175 0.267 0.354 720 0.264 0.228 0.233 0.219 0.280 0.635 0.219 0.318 0.426 Traffic 96 0.518 0.401 0.419 0.404 0.408 0.854 0.427 0.670 0.795 192 0.548 0.425 0.434 0.412 0.421 0.894 0.447 0.653 0.837 336 0.55 0.437 0.449 0.439 0.477 0.853 0.478 0.707 0.867 720 0.605 - - - - - - - - TTM Zero-shot Vs SOTA Few-shot (IMP) 1% 3% 2% 31% 6% 25% 32% TTM 5% Few-shot (IMP) 12% 14% 12% 38% 17% 32% 38% Table 1: Zero-shot and Few-shot 5% performance (MSE) of TTM and all SOTA models on D1 datasets for varying forecast lengths ($fl$).
 Bold and underscore denote the best and the second best.
@@
 SOTA few-shot results.
+IMP is calculated based on the mean of the percentage improvement of TTM achieved across all rows.
 Metric Data GPT4TS TTM (Ours) IMP (X) FT NPARAMS (M) Weather 3.2 0.29 (14X) Electricity 3.2 0.29 Traffic 5.5 0.29 TL NPARAMS (M) Weather 84.3 0.8 (106X) Electricity 84.3 0.8 Traffic 86.6 0.8 EPOCH TIME (s) Weather 37 1 (65X) Electricity 241 4 Traffic 782 8 MAX MEMORY (GB) Weather 21.9 1.06 (27X) Electricity 39.7 2.02 Traffic 53.7 1.34 TEST TIME (s) Weather 52.53 1.66 (54X) Electricity 353.4 8.07 Traffic 1163.9 13.5 Table 2: Computational Improvement in few-shot 10% for $fl$=96 using 1 A100 GPU.
@@
 4.1 Experimental Setting
-Datasets: Pre-training employs a subset of the Monash data hub Godahewa et al.
-(2021) of size $\sim 244$M samples.
+Datasets: Pre-training employs a subset of the Monash data hub [?] of size $\sim 244$M samples.
 We specifically exclude datasets (like yearly, monthly) as they do not possess sufficient length for the long-term forecasting task.
 Moreover, we remove all the datasets that we utilize for evaluation (i.e., weather, electricity, and traffic).
-For zero/few-shot evaluation we consider seven public datasets (D1): ETTH1, ETTH2, ETTM1, ETTM2, Weather, Electricity & Traffic as popularly used in most prior SOTA works Zhou et al.
-(2021); Nie et al.
-(2023).
-Since these datasets do not contain any exogenous variables nor exhibit cross-channel correlation benefits, we incorporate four other datasets (D2) for separately validating the efficacy of the decoder channel mixing and exogenous mixer module: bike sharing (BS) Fanaee-T (2013), carbon capture plant (CC) Jablonka et al.
-(2023), and 2 more datasets from Biz-IT observability domain ITB (2023): Application (APP) and Service (SER).
+For zero/few-shot evaluation we consider seven public datasets (D1): ETTH1, ETTH2, ETTM1, ETTM2, Weather, Electricity & Traffic as popularly used in most prior SOTA works [??].
+Since these datasets do not contain any exogenous variables nor exhibit cross-channel correlation benefits, we incorporate four other datasets (D2) for separately validating the efficacy of the decoder channel mixing and exogenous mixer module: bike sharing (BS) [?], carbon capture plant (CC) [?], and 2 more datasets from Biz-IT observability domain [?]: Application (APP) and Service (SER).
 Refer Appendix for full data details.
 SOTA Bechmarks: We benchmark TTM with the latest public SOTA forecasting models categorized as follows: [label=()] 1.
-LLM-based TS pre-trained models:GPT4TS Zhou et al.
-(2023), LLMTime Gruver et al.
-(2023) 2.
-Self-supervised pre-trained models: SimMTM Dong et al.
-(2023),Ti-MAE Li et al.
-(2023), TST Zerveas et al.
-(2021), LaST Wang et al.
-(2022), TF-C Zhang et al.
-(2022), CoST Woo et al.
-(2022) and Ts2Vec Yue et al.
-(2022) 3.
-TS transformer models:PatchTST Nie et al.
-(2023), FEDFormer Zhou et al.
-(2022), Autoformer Wu et al.
-(2021) 4.
-Other SOTA models:TSMixer Ekambaram et al.
-(2023), DLinear Zeng et al.
-(2022) and TimesNet Wu et al.
-(2022)
-Zero-shot Few-shot (10%) Data TTM (Ours) TTM (Ours) GPT4TS PatchTST TSMixer TimesNet Dlinear FEDFormer Autoformer ETTH1 0.428 0.422 0.590 0.633 0.578 0.869 0.691 0.639 0.702 ETTH2 0.362 0.360 0.397 0.415 0.370 0.479 0.605 0.466 0.488 ETTM1 0.545 0.458 0.464 0.501 0.441 0.677 0.411 0.722 0.802 ETTM2 0.302 0.278 0.293 0.296 0.282 0.320 0.316 0.463 1.342 Electricity 0.210 0.180 0.176 0.176 0.180 0.323 0.180 0.346 0.431 Traffic 0.555 0.428 0.440 0.430 0.429 0.951 0.446 0.663 0.749 Weather 0.236 0.227 0.238 0.242 0.236 0.279 0.241 0.284 0.300 TTM Zero-shot improvements -4% -2% -7% 27% 2% 27% 39% TTM Few-shot (10%) improvements 7% 9% 4% 34% 13% 34% 45% Table 3: Zero-shot and Few-shot 10% performance (MSE) of TTM over all SOTA models on D1 datasets, averaged across 4 forecast lengths.
+LLM-based TS pre-trained models:GPT4TS [?], LLMTime [?] 2.
+Self-supervised pre-trained models: SimMTM [?],Ti-MAE [?], TST [?], LaST [?], TF-C [?], CoST [?] and Ts2Vec [?] 3.
+TS transformer models:PatchTST [?], FEDFormer [?], Autoformer [?] 4.
+Other SOTA models:TSMixer [?], DLinear [?] and TimesNet [?]
+Zero-shot Few-shot (10%) Data TTM (Ours) TTM (Ours) GPT4TS PatchTST TSMixer TimesNet Dlinear FEDFormer Autoformer ETTH1 0.428 0.422 0.590 0.633 0.578 0.869 0.691 0.639 0.702 ETTH2 0.362 0.360 0.397 0.415 0.370 0.479 0.605 0.466 0.488 ETTM1 0.545 0.458 0.464 0.501 0.441 0.677 0.411 0.722 0.802 ETTM2 0.302 0.278 0.293 0.296 0.282 0.320 0.316 0.463 1.342 Electricity 0.210 0.180 0.176 0.176 0.180 0.323 0.180 0.346 0.431 Traffic 0.555 0.428 0.440 0.430 0.429 0.951 0.446 0.663 0.749 Weather 0.236 0.227 0.238 0.242 0.236 0.279 0.241 0.284 0.300 TTM Zero-shot Vs SOTA Few-shot (IMP) -4% -2% -7% 27% 2% 27% 39% TTM 10% Few-shot (IMP) 7% 9% 4% 34% 13% 34% 45% Table 3: Zero-shot and Few-shot 10% performance (MSE) of TTM over all SOTA models on D1 datasets, averaged across 4 forecast lengths.
 The numbers with bold and underscore denote the best and the second best results respectively.
+The last two rows show relative improvement (IMP) of TTM zero-shot and few-shot w.r.t.
+SOTA few-shot results.
+IMP is calculated based on the mean of the percentage improvement of TTM achieved across all rows.
 The full table is in Appendix.
@@
 Table 1 compares the performance of TTM model in zero-shot and few-shot (5%) settings across multiple $fl$s.
-Baseline results are reported from  Zhou et al.
-(2023) as we use the same few-shot data filtering strategy as followed in that paper.
+Baseline results are reported from  [?] as we use the same few-shot data filtering strategy as followed in that paper.
 Note that for zero-shot performance, the pre-trained TTM model is directly evaluated on the test set.
@@
 Since these models are based on LLaMA and are massive in size, the authors used only the last window of the standard test-set for faster evaluation, as opposed to using all the windows in the full test-set .
-Hence, we compare them separately in Table 4 based on the same datasets, $fl$, and test-set as reported in  Gruver et al.
-(2023).
+Hence, we compare them separately in Table 4 based on the same datasets, $fl$, and test-set as reported in  [?].
 In this evaluation, we outperform LLMTime by a substantial margin of 29%.
 In addition, there are alternative pre-training approaches like masked modeling and contrastive learning techniques that may not offer universal transfer learning capabilities across all datasets (like TTM), but they excel in enabling cross-transfer learning between two datasets when carefully selected.
-Table 2 illustrates the cross-transfer learning from ETTH2 to ETTH1 for these models in various few-shot settings (as reported in  Dong et al.
-(2023)).
+Table 2 illustrates the cross-transfer learning from ETTH2 to ETTH1 for these models in various few-shot settings (as reported in  [?]).
 Notably, TTM, with no specific cross-data selection, outperforms all popular SOTAs, including the latest SimMTM, by 17-43%.
@@
 Weather 0.321 0.236 0.267 0.233 0.264 0.231 0.240 0.229 0.238 0.227
-IMP 36% 21% 12% 9% 2%
+% IMP 36% 21% 12% 9% 2%
 Table 6: Improvement (IMP) of Pre-Training (PT) over Randomly Initialized (RI) weights. MSE averaged across all $fl$s.
@@
 Other SOTAs are not reported here, considering their inferior performance and space constraints.
-Specifically, we compare with TSMixer with channel-mixing enabled (TSMixer-CM) and TSMixer with Cross-Channel reconciliation head (TSMixer-CC) Ekambaram et al.
-(2023) as they are the latest SOTAs in channel-correlation modeling.
+Specifically, we compare with TSMixer with channel-mixing enabled (TSMixer-CM) and TSMixer with Cross-Channel reconciliation head (TSMixer-CC) [?] as they are the latest SOTAs in channel-correlation modeling.
 We can see that TTM-CM outperforms all the competitive models with a significant margin (15-44%), thus, demonstrating the power of TTM in capturing inter-channel correlations.
@@
 Thus, the performance impact of leveraging pre-trained weights greatly increases as the size of the training data and the available time-to-fine-tune reduces.
-Effect of Adaptive Patching: The Adaptive patching presented in Section 3.1.1 is a key component of the TTM backbone.
-Table 7 demonstrates the relative improvements when we employ adaptive patching with respect to a vanilla TTM backbone without this module.
+Effect of Adaptive Patching: The Adaptive patching presented in Section 3.1 is a key component of the TTM backbone.
+Table 4.2 demonstrates the relative improvements when we employ adaptive patching with respect to a vanilla TTM backbone without this module.
 We can see that adaptive patching adds an extra 4% improvement during zero-shot and 2% improvement for 10% few-shot.
@@
 Hence this component is optional and can be enabled when working with shorter input context length.
+4.6 Comparison with recent concurrent works
+In this section, we compare the performance of TTM with the very recent concurrent works published in 2024.
+Time-LLM [?] is a recent SOTA that enables successful reprogramming of LLMs for time-series tasks.
+However, as depicted in Table 10, TTM outperforms Time-LLM by 8% in Fewshot 5% and 2% in Fewshot 10% setting.
+Likewise, UniTime [?] proposes novel techniques towards language-empowered cross-domain time-series forecasting.
+As indicated in Table 11, TTM’s zero-shot results surpass UniTime’s zero-shot results by 27%.
+Further, TTM does not need any cross-domain transfer learning and pair-wise mapping as our pretraining is enabled on a diverse set of datasets.
+Data Few-shot (5%)) Few-shot (10%))
+TTM Time-LLM TTM Time-LLM
+ETTH1 0.393 0.627 0.422 0.556
+ETTH2 0.334 0.382 0.36 0.37
+ETTM1 0.44 0.425 0.458 0.404
+ETTM2 0.28 0.274 0.278 0.277
+Weather 0.231 0.26 0.227 0.234
+Electricity 0.179 0.179 0.18 0.175
+Traffic 0.421 0.423 0.428 0.429
+TTM IMP(%) 8% 2%
+Table 10: TTM Vs Time-LLM on D1 Datasets averaged across considered 4 forecast lengths. MSE reported.
+Data $fl$ Zero-shot
+UniTime TTM
+ETTH2 96 0.306 0.285
+192 0.389 0.341
+336 0.424 0.383
+720 0.434 0.441
+Electricity 96 0.409 0.169
+192 0.41 0.196
+336 0.439 0.209
+720 0.487 0.264
+Weather 96 0.21 0.154
+192 0.264 0.203
+336 0.326 0.256
+720 0.402 0.329
+TTM IMP (%) 27%
+Table 11: TTM vs UniTime
+MSE Improvement (IMP)
+in zero-shot setting.
 5 Conclusions and Future Work
@@
 Going forward, we plan to extend our approach to many other downstream tasks beyond forecasting for a purely foundational approach in time series.
+Appendix A
+The outline of our appendix section is as follows:
+  • Detailed Literature Survey - Section A.1
+  • TSMixer Background - Section A.2
+  • Details on Pre-training Datasets - Section A.3
+  • Details on Evaluation Datasets - Section A.4
+  • Baseline Implementation Details - Section A.5
+  • TTM Source Code and Implementation Details - Section A.6
+  • TTM Computational Benefits: Setup details - Section A.7
+  • Full Tables - Section A.8
+A.1 Detailed Literature Survey
+Multivariate Time Series Forecasting
+Statistical approaches for time series forecasting, such as SARIMAX and Exponential Smoothing, generally generate forecasts independently for each time series [?].
+These methods are essentially univariate and do not build a single model by learning from multiple time series.
+On the other hand, more advanced models, built upon machine/deep learning techniques, including LightGBM-based models [??], N-BEATS [?], and DeepAR [?], have the capability to learn from multiple time series.
+However, these models still follow univariate approaches, thus ignoring any potential cross-channel correlations.
+Advanced multivariate forecasting models mostly involve deep neural networks, specifically the transformer [?] architecture.
+A series of transformer-based model have been proposed in the last few years including Informer [?], Autoformer [?], and FEDFormer [?].
+Although these models outperformed all the prior arts, the DLinear [?] model showed that an embarrassingly simple linear model can beat these models by following a few empirically established steps like time series decomposition, normalization, and channel-independent modeling.
+PatchTST [?] showed that transformers can be effective for forecasting if the input time series is patched or segregated in multiple windows, and subsequently, modeled by a transformer.
+The patching operation helps preserve local semantic information, accommodates a longer history, and reduces computation time.
+The PatchTST model outperformed all prior transformer-based models and the DLinear model.
+Although PatchTST reinstated faith in transformers for time series modeling, transformer-based models are generally resource-intensive, with slow execution and a high memory footprint.
+The recently proposed TSMixer model [?] addresses these challenges effectively.
+TSMixer, built on the MLPMixer architecture [?], stands out for its exceptional speed and lightweight design.
+It has attained state-of-the-art (SOTA) performance on benchmark datasets, demonstrating a 2-3X reduction in both execution time and memory usage.
+Pre-trained Models for Time Series
+One major drawback of all the above models is that they need to be trained in-domain.
+Hence, none of these models can be transferred to out-of-domain data with zero or minimal training.
+This approach has been found to be extremely beneficial in the natural language processing (NLP) domain with the invention of BERT [?] and GPT [?] models.
+However, this is an extremely challenging task in the time series domain because of the unavailability of a publicly accessible large pre-training corpora.
+There are multiple independent time series datasets, but, unlike in NLP, these datasets differ significantly in important characteristics such as the domain of the data (e.g., retail, sensor data, traffic, etc.), the number of channels, temporal resolution, and length.
+This makes it hard to train a single model on all the datasets together.
+Hence, a few prior works have focused on experimenting with same-dataset self-supervised learning for time series [????].
+These methods learn a time series representation from the train split of a dataset, build a forecaster on top of the learned representation on the same data, and then evaluate it on the test split of the same dataset.
+Although these approaches have demonstrated promising results, they do not provide evidence of the transfer capability of the model between datasets.
+Recent works such as SimMTM [?] and TF-C [?] have demonstrated the transfer capabilities of their models between pairs of datasets.
+These pairs are carefully chosen so that the source (the dataset where the model is pre-trained) and target (the dataset where the model is fine-tuned and tested) datasets share some matching properties.
+For instance, SimMTM showcased its few-shot capability by selecting ETTH2 as the source data and ETTH1 as the target data.
+Both ETTH1 and ETTH2 are collected from Electricity Transformers at two stations, denoting data from a similar domain.
+TF-C demonstrated the transferability of the model across four different (source, target) pairs, such as (ECG, EMG) and (FD-A, FD-B), where domain-similarity exists in both the source and target datasets.
+Pre-trained LLMs for Time Series
+To tackle the aforementioned challenges, there has been a notable increase in the adoption of pre-trained large language models (LLMs) for time series tasks.
+These models are approached as cross-domain transfer learning problems.
+The LLMTime model [?] feeds the time series values as text representations and demonstrates promising performance in a zero-shot setting.
+The GPT4TS model [?] adopts a pre-trained LLM like GPT and fine-tunes only the input embedding layer, normalization layers, and output layer.
+Specifically, it does not alter the self-attention weights and feed-forward layers.
+This approach to building a pre-trained model for time series from LLMs is promising, but it does not model cross-channel correlations observed in many multivariate time series datasets.
+Moreover, these LLMs are very large and exhibit slow execution and a large memory footprint.
+Dataset Resolution
+solar_10_minutes_dataset.tsf 10_minutes
+australian_electricity_demand_dataset.tsf half_hourly
+solar_4_seconds_dataset.tsf 4_seconds
+wind_4_seconds_dataset.tsf 4_seconds
+us_births_dataset.tsf daily
+saugeenday_dataset.tsf daily
+sunspot_dataset_without_missing_values.tsf daily
+australian_weather_dataset.tsf daily
+kdd_cup_2018_dataset_without_missing_values.tsf hourly
+bitcoin_dataset_without_missing_values.tsf daily
+wind_farms_minutely_dataset_without_missing_values.tsf minutely
+australian_electricity_demand_dataset_downsample_2.tsf hourly
+solar_4_seconds_dataset_downsample_225.tsf 15_minutes
+wind_4_seconds_dataset_downsample_900.tsf hourly
+london_smart_meters_dataset_without_missing_values_downsample_2.tsf hourly
+solar_4_seconds_dataset_downsample_900.tsf hourly
+wind_farms_minutely_dataset_without_missing_values_downsample_10.tsf 10_minutes
+solar_10_minutes_dataset_downsample_6.tsf hourly
+wind_4_seconds_dataset_downsample_150.tsf 10_minutes
+wind_farms_minutely_dataset_without_missing_values_downsample_15.tsf 15_minutes
+solar_4_seconds_dataset_downsample_150.tsf 10_minutes
+wind_4_seconds_dataset_downsample_225.tsf 15_minutes
+wind_farms_minutely_dataset_without_missing_values_downsample_60.tsf hourly
+Table 12: List of pre-training datasets.
+Datasets with “_downsample_X.tsf” suffix denotes an augmented dataset created from the original dataset by downsampling the latter.
+Please note that, these pre-training datasets have no overlap with the evaluation datasets.
+Specifically, the australian_electricity_demand_dataset and australian_weather_dataset used in pre-training are completely different (w.r.t location, measured variables, type, resolution, length, etc.) from the standard Electricity (ECL) and Weather dataset used in the evaluation.
+A.2 TSMixer Background
+We employed TSMixer [?] as a building block for the proposed TTM model due to its state-of-the-art performance, faster execution, and significantly lower memory usage.
+However, as explained in the main paper, vanilla TSMixer cannot be trained on multiple diverse datasets.
+Therefore, it necessitated the incorporation of the proposed novel components.
+In this section, we provide a high-level overview of the TSMixer model for a simpler and quicker understanding by the readers.
+TSMixer is a lightweight alternative to transformer-based time series models, with no compromise on forecast accuracy.
+TSMixer adopts some well-established pre-processing steps from the literature, such as normalization and patching.
+Additionally, it offers the flexibility of enabling or disabling channel mixing.
+Channel mixing has been found to be beneficial in handling multivariate datasets with cross-channel correlations.
+For the main learning process, TSMixer employs a series of MLPMixer [?] blocks that perform inter-patch, intra-patch, and inter-channel mixing operations.
+A mixing operation in TSMixer ensures learning correlations across a specific dimension.
+For example, inter-channel mixing enables it to learn cross-channel correlations.
+In the experiments, we employed three different flavors of the TSMixer model: TSMixer vanilla, TSMixer with cross-channel mixing enabled (TSMixer-CM), and TSMixer with cross-channel reconciliation head (TSMixer-CC).
+We request the authors to refer to [?] for further details about these variants.
+A.3 List of Pre-training Datasets
+We employ a subset of the datasets available in the Monash forecasting data repository [?] available at https://forecastingdata.org/.
+Since our primary focus in this study is long term forecasting with forecast length ranging from 96 to 720, it is not possible to use yearly, monthly, quarterly, or weekly datasets due to their short lengths.
+Hence, we skip a few datasets of short lengths.
+The final list of all pre-training datasets is shown in Table 12.
+Temporal cross validation [?] is used to chronologically split all the time series into train and validation parts.
+During pre-training, moving windowing technique is used to create $\left(\bm{X},\bm{Y}\right)$ pairs of lengths $sl$ and $fl$ respectively.
+During pre-training, the total number of train and validation samples (i.e., number of $\left(\bm{X},\bm{Y}\right)$ pairs) are $244$M and $71$M respectively.
+Please note that, these pre-training datasets have no overlap with the evaluation datasets.
+In specific, the australian_electricity_demand_dataset and australian_weather_dataset used in pre-training are completely different (w.r.t location, measured variables, type, resolution, length, etc.) from the standard Electricity (ECL) and Weather dataset used in the evaluation.
+Set Dataset Resolution Length Total #Channels #Target variables #Exog. variables
+Source
+D1 ETTH1 1 hour 17420 7 7 Not Applicable https://github.com/zhouhaoyi/ETDataset/tree/main/ETT-small
+ETTH2 1 hour 17420
+ETTM1 15 minutes 69680
+ETTM2 15 minutes 69680
+Weather 10 minutes 52696 21 21
+https://www.bgc-jena.mpg.de/wetter/
+ECL 1 hour 26304 321 321
+https://archive.ics.uci.edu/dataset/321/electricityloaddiagrams20112014
+Traffic 1 hour 17544 862 862
+https://pems.dot.ca.gov/
+D2 BS 1 hour 17379 14 3 11
+https://archive.ics.uci.edu/dataset/275/bike+sharing+dataset
+CC 2 minutes 5409 8 2 5
+https://github.com/kjappelbaum/aeml/blob/main/paper/20220210_smooth_window_16.pkl
+APP 10 minutes 8834 39 4 35
+https://github.com/BizITObs/BizITObservabilityData/tree/main/Complete/Time%20Series/RobotShop
+SER 10 minutes 8835 107 72 35
+https://github.com/BizITObs/BizITObservabilityData/tree/main/Complete/Time%20Series/RobotShop
+Table 13: Details of the evaluation datasets.
+A.4 List of Evaluation Datasets
+Table 13 illustrates various characteristics of the eleven evaluation datasets. Below, we present the details.
+Set D1
+For zero/few/full-shot evaluation, we utilize seven multivariate time series datasets that have consistently been employed in the literature. Below, we offer a brief overview of these datasets.
+  1. 1.
+    ETT datasets: The four ETT datasets [?] (ETTH1, ETTH2, ETTM1, ETTM2) contain multivariate time series data collected from electrical transformers at two stations.
+ETTH1 and ETTH2 are collected at an hourly interval, while ETTM1 and ETTM2 are collected every 15 minutes.
+All four datasets have 7 channels.
+  2. 2.
+    Weather: The weather dataset consists of 21 channels, which serve as weather indicators. It is collected at 10-minute intervals at the Max Planck Institute of Biogeochemistry weather station.
+  3. 3.
+    Electricity (ECL): The Electricity dataset, also known as the ECL dataset, comprises the hourly electricity consumption data of 321 clients.
+  4. 4.
+    Traffic: This dataset records the hourly rates of road occupancy on the San Francisco Freeways using 862 sensors.
+We used the datasets provided in the repository of the Autoformer paper [?] [^5] Available at https://github.com/thuml/Autoformer.
+For all the D1 datasets, we execute the same train/validation/test splitting as was performed in the literature [????].
+Set D2
+To assess the effectiveness of the proposed TTM model in extracting information from exogenous channels, we conduct evaluations on four additional datasets that are known to contain exogenous or control variables.
+  1. 1.
+    Bike Sharing (BS): The Bike Sharing dataset [?] documents the hourly rental counts of bikes from the Capital Bikeshare system in Washington D.C., USA, spanning the years 2011 to 2012.
+Rental counts are typically associated with environmental and seasonal conditions.
+Consequently, this 14-channel dataset encompasses various weather-related features.
+Our goal was to forecast all three rental counts: “casual”, “registered”, and “cnt” (total count).
+As the remaining 11 features are consistently available at all future time points, they are treated as exogenous variables in our experiment.
+  2. 2.
+    Carbon Capture Plant (CC): The Carbon Capture Plant data [?] records the emission profiles of “2-amino-2-methyl-1-propanol” (AMP) and “piperazine” (Pz) collected at every 2 minutes interval.
+We utilize the 8-channel dataset made available in the official repository of [?].
+Among the remaining 6 channels, the following 5 serve as control variables: [“TI-19”,“FI-19”, “TI-3”, “FI-11”, “TI-1213”].
+The remaining 1 variable is treated as a conditional variable (as it is neither a target variable nor available during the forecast period to consider it as exogenous).
+For additional details, please refer to the supplementary materials of [?].
+  3. 3.
+    Service (SER): This dataset pertains to the cloud-based “Stan’s Robot Shop” application, managed by Instana.
+It simulates a user’s e-commerce experience, encompassing site access to shipping, utilizing a load generator.
+Intermittent fault injection introduces diverse IT events.
+The dataset provides business KPIs for services (e.g., payment, catalog) and IT events tracked by Instana.
+Sampling occurs every 10 seconds due to high traffic and event frequency.
+For our experiments, all business KPIs are treated as target variables and IT events are treated as exogenous variables and the goal of our forecasting is to predict the business KPIs given the IT events.
+  4. 4.
+    Application (APP): This dataset is similar to the SER data, but it captures KPIs for the entire application instead of capturing at the service level.
+Even in this case, all business KPIs are treated as target variables and IT events are treated as exogenous variables and the goal of our forecasting is to predict the business KPIs given the IT events.
+A.5 Baseline Implementation Details
+We report the implementation details for all the baselines in Table 14.
+Category Baseline Used in Table Results Generated From Link to the used implementation
+(a) LLM- based pre-trained TS models GPT4TS Few-shot 5% in Table 1 Result from [?] Table 12
+N/A
+GPT4TS Runtime: Table 2 Ran Official Implementation
+GPT4TS Repository: https://github.com/DAMO-DI-ML/NeurIPS2023-One-Fits-All
+GPT4TS Few-shot 10% in Table 3 Result from [?] Table 13
+N/A
+GPT4TS Exogeneous in Table 5 Ran Official Implementation
+GPT4TS Repository: https://github.com/DAMO-DI-ML/NeurIPS2023-One-Fits-All
+LLMTime Table 4 Result from LLMTime Repository
+LLMTime Repository: https://github.com/ngruver/llmtime
+Result File: https://github.com/ngruver/llmtime/blob/main/precomputed_outputs/deterministic_csvs/autoformer_comb_results.csv
+(b) Self- supervised pre-trained models SimMTM Table 2 Directly reported from SimMTM paper N/A
+Ti-MAE
+TST
+LaST
+TF-C
+CoST
+TS2Vec
+(c) TS transformer models PatchTST Few-shot 5% and 10% in Table 1 and 3 Result from [?] Table 12, 13
+PatchTST Repository: https://github.com/yuqinie98/PatchTST
+FEDFormer
+N/A
+Autoformer
+N/A
+(d) Other SOTA Models TSMixer Used Official Implementation
+TSMixer Implementation: https://huggingface.co/docs/transformers/model_doc/patchtsmixer
+DLinear Result from [?] Table 12, 13
+N/A
+TimesNet Result from [?] Table 12, 13
+N/A
+(e) Concurrent Works Time-LLM Table 10 Result from [?] Table 3,4
+N/A
+UniTime Table 11 Results from [?] Table 5
+N/A
+Table 14: Implementation details for the baseline algorithms.
+A.6 TTM Source Code and Implementation Details
+Pre-training
+For our experiments, we build 5 pre-trained models using the Monash datasets for the following $(sl$,$fl)$ configurations: (512,96), (512,192), (512, 336), (512,720) and (96,24).
+Since our pre-trained models are very small, they can be trained quickly in a few hours (4-8 hrs based on $fl$ length), as opposed to several days/weeks in standard approaches.
+Hence, pre-training multiple TTM models is no longer a practical constraint.
+Pre-training is performed in a distributed fashion with 50 CPUs and 6 NVIDIA A100 GPUs with 40 GB GPU memory.
+Based on the $sl$ and $fl$ requirement of the target dataset, a suitable pre-trained model is selected to bootstrap the weights accordingly for fine-tuning.
+Other model configurations are as follows: patch length $pl$ = 64 (when $sl$ is 512) and 8 (when $sl$ is 96), stride $s$ = $pl$ (i.e.
+non-overlapping patches), number of patches $n=sl/pl$, number of levels in backbone $L$ = 6, number of TTM blocks per level $M$ = 2, number of decoder layers = 2, batch size $b$ = 3000, number of epochs $ep$ = 20, and dropout $do$ = 0.2.
+TSMixer-specific hyperparameters include feature scaler $fs$ = 3, hidden feature size $hf$ = $fs*pl$, expansion feature size $ef=hf*2$.
+Please note that $hf$ and $n$ will change across TTM blocks based on the adaptive patching strategy.
+Resolution prefix tuning is disabled by default and enabled only for shorter context lengths (as explained in Table.
+9).
+Decoder channel-mixing and exogenous mixer blocks are disabled during pre-training.
+Fine-tuning
+We have 2 sets of target datasets: D1 and D2 on which we fine-tune and test our performance.
+All D1 datasets use $sl$ = 512 and $fl$ is varied across {96, 192, 336, 720} for zero/few-shot forecasting.
+Likewise, in D2 datasets, BS data use $sl$ = 512 and $fl$ = 96 and remaining datasets use $sl$ = 96, $fl$ = 24.
+In addition, head dropout is set to 0.7 for smaller ETT datasets and 0.2 for other datasets.
+Likewise, the batch size is set to 8 for Traffic, 32 for Electricity, and 64 for all other datasets.
+All other parameters remain the same as used in the pre-training.
+Also, Decoder channel-mixing and exogenous mixer block are enabled for the D2 datasets.
+Unlike pre-training, fine-tuning is executed in just 1 A100 GPU as it is a fast process.
+All these hyperparameters are selected based on the validation performance, and the final test results are reported in the paper.
+Source Code
+For more implementation details on the TTM model, please refer to the source code of the important classes used in TTM.
+We have anonymized the important Python class files and shared them in the same technical appendix zip file.
+Important modules are listed below.
+The full project with complete reproducibility will be open-sourced on GitHub after the double-blind review.
+  • Class TinyTimeMixerConfig defines the requied configuration.
+  • Class TinyTimeMixerBlock implements the basic TSMixer Block
+  • Class ForecastChannelHeadMixer implements Exogenous Mixer Block
+  • Class TinyTimeMixerAdaptivePatchingBlock implements the Adaptive Patching statergy
+  • Class TinyTimeMixerDecoder implements the TTM Decoder
+  • Class TinyTimeMixerForPredictionHead implements the Forecasting Head
+  • Class TinyTimeMixerPreTrainedModel implements the pretraining model interfaces.
+  • Class TinyTimeMixerPatchify implements the required patching.
+  • Class TinyTimeMixerEncoder implements the TTM Backbone
+  • Class TinyTimeMixerModel implements the TTM Model wrappers.
+A.7 Computational Benefits of TTM over GPT4TS: Setup details
+Table  2 compares the computational benefits of TTM over GPT4TS, the most popular LLM-TS model and the current best SOTA in a few-shot setting.
+This section explains the experimental setup followed to enable this comparison.
+To execute GPT4TS with the best parameters, we used their official implementation as mentioned in Table 14.
+For a fair comparison, we run both models with the best-reported parameters in a single A100 GPU environment.
+Multi-GPU is avoided in this experiment to avoid IPC overheads for precise metric measurements.
+Since GPT4TS processes data in a purely univariate fashion while TTM fine-tuning processes data in a multi-variate fashion, we set the batch size accordingly to ensure that the number of univariate samples processed in each batch is the same for both models.
+For example, in TTM, if the batch size is set to 64, it implies processing 64 multivariate time series in a batch.
+Consequently, the equivalent batch size for GPT4TS (which process only univariate samples) is $64\times c$, where $c$ represents the number of channels in the dataset.
+Additionally, due to GPT4TS encountering out-of-memory (OOM) errors with the default high batch sizes used in TTM, we employed reduced batch sizes for this experiment.
+The following batch sizes were utilized for TTM: Weather (64), Electricity (8), and Traffic (2).
+The corresponding batch sizes for GPT4TS can be calculated by multiplying the TTM batch size with the channel counts of the respective dataset.
+Since Traffic and Electricity have very high number of channels, we had to significantly reduce its batch size for a consistent comparison across models.
+A.8 Full tables
+Here, we present the complete versions of various tables in the main paper.
+These full versions essentially include the test results for multiple forecast lengths ($fl$) across all datasets.
+Occasionally, these results are averaged across forecast lengths to conserve space in the main paper.
+Full table for 10% few-shot experiment
+Table 15 shows the 10% few-shot results for all forecast lengths across all D1 datasets.
+Full table for validating effect of pre-training
+Table 16 shows the effect of pre-training when compared to random initialization of the model weights across all D1 datasets for all forecast lengths.
+Full table for validating adaptive patching
+Table 17 provides a comprehensive overview, systematically validating the impact of adaptive patching across all D1 datasets and forecast lengths.
+Full table for validating effect of downsampling
+Table 18 offers a comprehensive summary, systematically validating the influence of dataset augmentation through downsampling across all D1 datasets and forecast lengths (96 and 192).
+Zero-shot Few-shot (10%)
+FL TTM (Ours) TTM (Ours) GPT4TS PatchTST TSMixer TimesNet Dlinear FEDFormer Autoformer
+ETTH1 96 0.365 0.374 0.458 0.516 0.472 0.861 0.492 0.512 0.613
+192 0.393 0.388 0.570 0.598 0.538 0.797 0.565 0.624 0.722
+336 0.415 0.395 0.608 0.657 0.604 0.941 0.721 0.691 0.750
+720 0.538 0.533 0.725 0.762 0.698 0.877 0.986 0.728 0.721
+ETTH2 96 0.285 0.284 0.331 0.353 0.288 0.378 0.357 0.382 0.413
+192 0.341 0.335 0.402 0.403 0.370 0.490 0.569 0.478 0.474
+336 0.383 0.383 0.406 0.426 0.400 0.537 0.671 0.504 0.547
+720 0.441 0.439 0.449 0.477 0.422 0.510 0.824 0.499 0.516
+ETTM1 96 0.413 0.377 0.390 0.410 0.373 0.583 0.352 0.578 0.774
+192 0.476 0.424 0.429 0.437 0.413 0.630 0.382 0.617 0.754
+336 0.553 0.424 0.469 0.476 0.451 0.725 0.419 0.998 0.869
+720 0.737 0.606 0.569 0.681 0.527 0.769 0.490 0.693 0.810
+ETTM2 96 0.187 0.173 0.188 0.191 0.185 0.212 0.213 0.291 0.352
+192 0.261 0.237 0.251 0.252 0.251 0.270 0.278 0.307 0.694
+336 0.323 0.303 0.307 0.306 0.302 0.323 0.338 0.543 2.408
+720 0.436 0.400 0.426 0.433 0.392 0.474 0.436 0.712 1.913
+Weather 96 0.154 0.148 0.163 0.165 0.159 0.184 0.171 0.188 0.221
+192 0.203 0.194 0.210 0.210 0.206 0.245 0.215 0.250 0.270
+336 0.256 0.245 0.256 0.259 0.256 0.305 0.258 0.312 0.320
+720 0.329 0.320 0.321 0.332 0.321 0.381 0.320 0.387 0.390
+Electricity 96 0.169 0.139 0.139 0.140 0.141 0.299 0.150 0.231 0.261
+192 0.196 0.157 0.156 0.160 0.158 0.305 0.164 0.261 0.338
+336 0.209 0.179 0.175 0.180 0.182 0.319 0.181 0.360 0.410
+720 0.264 0.244 0.233 0.223 0.238 0.369 0.223 0.530 0.715
+Traffic 96 0.518 0.398 0.414 0.403 0.404 0.719 0.419 0.639 0.672
+192 0.548 0.415 0.426 0.415 0.419 0.748 0.434 0.637 0.727
+336 0.550 0.422 0.434 0.426 0.425 0.853 0.449 0.655 0.749
+720 0.605 0.477 0.487 0.474 0.469 1.485 0.484 0.722 0.847
+Table 15: Zero-shot and Few-shot 10% performance (MSE) on TTM and all SOTA models on seven datasets for varying forecast lengths ($fl$).
+The numbers with bold and underscore denote the best the second best results respectively.
+FL Zero-shot Few-shot (5%)) Few-shot (10%))
+5 Epochs 50 Epochs 5 Epochs 50 Epochs
+RI PT RI PT RI PT RI PT RI PT
+ETTH1 96 0.737 0.365 0.671 0.366 0.548 0.366 0.599 0.372 0.434 0.374
+192 0.737 0.393 0.682 0.392 0.666 0.391 0.626 0.388 0.492 0.388
+336 0.731 0.415 0.708 0.408 0.674 0.421 0.621 0.399 0.588 0.395
+720 0.735 0.538 - - - - 0.716 0.516 0.695 0.533
+ETTH2 96 0.380 0.285 0.354 0.283 0.342 0.282 0.293 0.284 0.287 0.284
+192 0.392 0.341 0.391 0.337 0.404 0.338 0.374 0.331 0.368 0.335
+336 0.394 0.383 0.390 0.382 0.390 0.383 0.390 0.380 0.399 0.383
+720 0.432 0.441 - - - - 0.428 0.434 0.424 0.439
+ETTM1 96 0.707 0.413 0.367 0.384 0.345 0.359 0.352 0.388 0.354 0.377
+192 0.716 0.476 0.446 0.417 0.378 0.402 0.401 0.408 0.378 0.424
+336 0.726 0.553 0.440 0.463 0.423 0.424 0.425 0.437 0.408 0.424
+720 0.737 0.737 0.641 0.668 0.645 0.575 0.505 0.598 0.471 0.606
+ETTM2 96 0.296 0.187 0.203 0.175 0.190 0.174 0.187 0.174 0.178 0.173
+192 0.325 0.261 0.258 0.237 0.247 0.240 0.244 0.231 0.236 0.237
+336 0.360 0.323 0.313 0.295 0.309 0.299 0.297 0.292 0.289 0.303
+720 0.430 0.436 0.411 0.391 0.412 0.407 0.390 0.388 0.392 0.400
+Weather 96 0.277 0.154 0.181 0.153 0.179 0.152 0.165 0.150 0.161 0.148
+192 0.302 0.203 0.241 0.201 0.229 0.198 0.209 0.192 0.210 0.194
+336 0.331 0.256 0.280 0.253 0.283 0.250 0.262 0.252 0.258 0.245
+720 0.375 0.329 0.367 0.325 0.367 0.326 0.322 0.321 0.324 0.32
+Electricity 96 0.874 0.169 0.212 0.147 0.148 0.142 0.181 0.144 0.141 0.139
+192 0.874 0.196 0.260 0.174 0.174 0.162 0.192 0.164 0.157 0.157
+336 0.893 0.209 0.273 0.187 0.192 0.184 0.225 0.183 0.179 0.179
+720 0.917 0.264 0.678 0.245 0.282 0.228 0.281 0.251 0.237 0.244
+Traffic 96 1.434 0.518 0.504 0.421 0.413 0.401 0.432 0.405 0.402 0.398
+192 1.438 0.548 0.539 0.457 0.428 0.425 0.495 0.423 0.420 0.415
+336 1.460 0.550 0.736 0.485 0.482 0.437 0.488 0.435 0.426 0.422
+720 1.481 0.605 - - - - 0.531 0.520 0.471 0.477
+Table 16: Effect of pre-training (PT) when compared to random initialization (RI) of model weights. MSE reported.
+Data FL Zero-shot Few-shot (10%)
+Vanilla Adaptive Patching Vanilla Adaptive Patching
+ETTH1 96 0.369 0.365 0.367 0.374
+192 0.395 0.393 0.387 0.388
+336 0.457 0.415 0.408 0.395
+720 0.784 0.538 0.751 0.533
+ETTH2 96 0.283 0.285 0.282 0.284
+192 0.346 0.341 0.345 0.335
+336 0.400 0.383 0.393 0.383
+720 0.470 0.441 0.469 0.439
+ETTM1 96 0.446 0.413 0.38 0.377
+192 0.498 0.476 0.413 0.424
+336 0.598 0.553 0.515 0.424
+720 0.765 0.737 0.544 0.606
+ETTM2 96 0.191 0.187 0.171 0.173
+192 0.268 0.261 0.232 0.237
+336 0.333 0.323 0.284 0.303
+720 0.447 0.436 0.400 0.400
+Weather 96 0.159 0.154 0.149 0.148
+192 0.204 0.203 0.196 0.194
+336 0.261 0.256 0.249 0.245
+720 0.333 0.329 0.328 0.32
+Electricity 96 0.179 0.169 0.138 0.139
+192 0.188 0.196 0.157 0.157
+336 0.220 0.209 0.181 0.179
+720 0.274 0.264 0.238 0.244
+Traffic 96 0.521 0.518 0.410 0.398
+192 0.536 0.548 0.416 0.415
+336 0.559 0.550 0.422 0.422
+720 0.609 0.605 0.475 0.477
+Table 17: Effect of adapting patching. MSE reported.
+Data FL Zero-shot
+Original data With Downsampling
+ETTH1 96 0.501 0.365
+192 0.706 0.393
+ETTH2 96 0.309 0.285
+192 0.392 0.341
+ETTM1 96 0.829 0.413
+192 0.876 0.476
+ETTM2 96 0.222 0.187
+192 0.300 0.261
+Weather 96 0.163 0.154
+192 0.208 0.203
+Electricity 96 0.334 0.169
+192 0.514 0.196
+Traffic 96 0.838 0.518
+192 0.958 0.548
+Table 18: Effect of downsampling. MSE reported.